### v1

# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

Abstract.
Containerization in a multi-cloud environment facilitates workload portability and optimized resource utilization. Containerization in multi-cloud environments has received significant attention in recent years both from academic research and industrial development perspectives. However, there exists no effort to systematically investigate the state of research on this topic. The aim of this research is to systematically identify and categorize the multiple aspects of containerization in multi-cloud environment. We conduct the Systematic Mapping Study (SMS) on the literature published between January 2013 and March 2023. Eighty-six studies were finally selected and the key results are: (1) Four leading themes on cloud computing and network systems research were identified: ‘Scalability and High Availability’, ‘Performance and Optimization’, ‘Security and Privacy’, and ‘Multi-Cloud Container Monitoring and Adaptation’. (2) Seventy-four patterns and strategies for containerization in multi-cloud environment were classified across 10 subcategories and 4 categories. (3) Ten quality attributes considered were identified with 47 associated tactics. (4) Four frameworks were introduced from the analysis of identified challenges and solutions: security, automation, deployment, and monitoring challenge-solution frameworks. The results of this SMS will assist researchers and practitioners in pursuing further studies on containerization in multi-cloud environment and developing specialized solutions.
Containerization, Multi-Cloud Environment, Challenges, Solutions, Frameworks, Systematic Mapping Study
journal: TOSEM
1. Introduction
The use of containers in multi-cloud environment has been widespread in the industry for many years (Naik, 2016b). Containers are standalone and executable packages of software that include everything needed to run an application, such as code, system tools, libraries, and settings (Merkel et al., 2014). Containers allow the packaging of an application along with its required dependencies, which makes it easy to move the application across different environments with little or no modification (Docker, 2021). On the other hand, multi-cloud environment is the distribution of cloud assets (Petcu, 2013). Using containers in multi-cloud environment allows organizations to achieve flexibility, agility, and cost efficiency. Developers can build applications in a consistent environment and easily move them between multiple cloud platforms, thereby leveraging the unique strengths (e.g., extensive infrastructure, seamless integration, advanced data analytic) of each platform.
Containerization in multi-cloud is illustrated in Figure 1. This figure provides an overview of applications deployed within various cloud configurations including public, private, and hybrid models. Each cloud hosts multiple applications encapsulated in containers, and completes different types of tasks with their required binaries and libraries, emphasizing the portability and isolation that containerization offers. A container platform layer, which could be exemplified by systems like Kubernetes or Docker, manages and orchestrates these containers across the different cloud environments. This architecture enables a flexible and scalable approach, allowing efficient application deployment and operations in a multi-cloud setting, while maintaining consistency and resilience across platforms.
Figure 1. Context: Containerization in multi-cloud environment
However, utilizing containerization in multi-cloud environment is not without challenges that organizations may face (Saif et al., 2022). These challenges may arise during different phases and activities of container-based application development including architectural design phase, system implementation, or when establishing an automated development infrastructure. Furthermore, the challenges can also present themselves during system testing, the coding process, and the deployment phase. For instance, during the architectural design phase, challenges may involve ensuring seamless integration of containerized components with existing systems and selecting appropriate container orchestration patterns, strategies, and tools that can effectively manage containers across multiple clouds (Raj et al., 2018) (Baby and Vysala, 2015), and while establishing an automated development infrastructure, organizations may encounter difficulties in creating efficient Continuous Integration/Continuous Deployment (CI/CD) pipelines that handle multi-cloud deployment scenarios while maintaining consistent performance and security standards (Debroy et al., 2018). Similarly, In the system implementation phase, challenges may include dealing with potential compatibility issues between containers and various cloud platforms, as well as optimizing resource utilization to control costs effectively (Helali and Omri, 2021).
Recent research (e.g., see Table 8) has highlighted the growing significance of container utilization in multi-cloud environment. These studies have explored various aspects, including container roles and strategies, architectural patterns, Quality Attributes (QAs), and the tools and frameworks employed. Additionally, these studies have explored the challenges associated with automation, deployment, monitoring, and security of containerization applications in multi-cloud environment. Despite the breadth of knowledge available, such valuable information is dispersed across different publications, spanning scientific research papers and the grey literature. This fragmentation of knowledge about containerization applications in multi-cloud environment presents a significant navigational challenge for practitioners tasked with real-world implementations. They need to thoroughly scrutinize numerous aspects, such as patterns and strategies for containerization applications in multi-cloud environment, to locate specific information relevant to their use cases, whether it be a pattern, a challenge, or a solution to an existing problem. Consequently, this spreading of knowledge restricts the development of a holistic understanding of the containerization applications in multi-cloud environment, leaving practitioners underprepared to devise comprehensive solutions.
Furthermore, this state of affairs of containerization in multi-cloud also poses challenges for the academic community. Researchers bear the obligation of navigating through this vast array of information to uncover the precise aspects they seek - be it a unique pattern, an unexplored challenge, or an innovative solution. Moreover, the distributed nature of this knowledge delays the formation of a cohesive and unified understanding of the subject matter. Recent studies (e.g., (Bordeleau et al., 2020) (Mihai et al., 2022) (Sarker, 2022)) have notably highlighted that challenges in design, development, monitoring, and testing of containerization applications in multi-cloud environment are deeply interconnected with the software development life cycle.
To address this knowledge gap, we established the objective of this research is to systematically identify and categorize various facets of container utilization in multi-cloud environment. These facets encompass the roles of containers, their implementation strategies, architectural patterns, quality attributes, as well as the motivations of the customers. Additionally, we aspire to explore the challenges and solutions related to automation, deployment, monitoring, and security of containerization applications in multi-cloud environment. To achieve this objective, we employed a systematic approach to conduct a mapping study. We finally selected 86 studies that focus on containerized applications in multi-cloud environment for further analysis.
The empirical findings and frameworks of this Systematic Mapping Study (SMS) offer significant potential benefits for developers of container-based applications. By exploring a diverse range of implementation strategies, architectural patterns, and quality attributes, practitioners can gain valuable insights to enhance their practices. Moreover, a detailed examination of challenges and solutions related to automation, deployment, monitoring, and security further provides practitioners with the knowledge to address problems effectively and improve dedicated quality attributes. Additionally, these findings and frameworks offer valuable support to practitioners in comprehending various patterns, strategies, and tactics specific to containerized applications in multi-cloud environment. The proposed frameworks can act as a facilitator of knowledge sharing, training, and ensure a shared understanding among practitioners regarding the complexities and solutions within containerized applications in multi-cloud environment. Ultimately, this SMS serves to empower software developers and practitioners, enabling them to optimize their approaches and elevate the quality and efficiency of containerized applications in multi-cloud environment.
Our SMS makes the following key contributions:
  • We developed four frameworks to address challenges and solutions in security, automation, deployment, and monitoring for containerization in multi-cloud environment.
  • We systematically identified and classified an extensive list of patterns and strategies for container-based applications in multi-cloud environment.
  • We systematically identified and classified quality attributes and tactics for containerized applications in multi-cloud environment.
  • We identified and classified a wide range of tools and frameworks for containerized applications in multi-cloud environment.
  • We have publicly released a dataset, which is available online (Waseem et al., 2024), to enable researchers and practitioners to access all collected data, replicate, and validate our SMS. This dataset includes detailed hierarchies of the developed frameworks, identified and classified container implementation strategies, and the roles of containers in multi-cloud environment. It also encompasses patterns, strategies, quality attributes, tactics, tools, and frameworks for containerized-based applications in multi-cloud environment.
The paper is structured as follows: Section 2 describes the research methodology employed. Section 3 presents the results and findings of this SMS. Section 4 discusses the implications of the SMS results. Section 5 clarifies the potential threats to the validity of this SMS. Section 6 reviews relevant prior work and highlights the differences between our work and previous research. Section 7 draws conclusions and outlines future research directions.
2. Methodology
We conducted a Systematic Mapping Study (SMS) by following the guidelines in (Petersen et al., 2008a) and augmenting them with SLR strategies (Keele et al., 2007). Our SMS consists of three phases: specifying research questions and search string, conducting the literature search, and performing data analysis and documentation. Figure 2 illustrates the SMS execution process.
Figure 2. Schematic representation of the research methodology implemented in this study
2.1. Research Questions
To conduct this SMS, we formulated six Research Questions (RQs), as outlined in Table 1, aligning with the objective of our SMS. The table provides a structured view of the RQs, their corresponding rationale, and categorization.
Table 1. Research questions and their rationale
# Research Questions Rationale
Container Roles and Implementation
RQ1
What roles do containers play, and what strategies can be employed for implementing container-based applications in multi-cloud environment?
To investigate the role of containers and explore implementation strategies in multi-cloud environment, with the objective of enhancing understanding and facilitating efficient deployment and management of applications across multiple cloud environments.
RQ2
What architectural patterns and strategies are utilized in the implementation of container-based applications in multi-cloud environment?
To explore and document the various architectural patterns and strategies that are utilized in implementing container-based applications in multi-cloud environment, thus providing a roadmap for efficient and scalable application development and deployment.
RQ3
What are the quality attributes and the tactics associated with their implementation in container-based applications in multi-cloud environment?
To examine and analyze the quality attributes required for container-based applications in multi-cloud environment and the corresponding tactics employed in their implementation, enabling the identification of best practices and guidelines for developing robust and reliable applications in such environments.
Challenges and Solutions
RQ4
What are the challenges related to security, automation, deployment, and monitoring in container-based applications in multi-cloud environment?
To comprehensively evaluate and gain insights into the challenges surrounding security, automation, deployment, and monitoring concerning container-based applications in multi-cloud environment, this question aims to provide valuable insights into the obstacles and intricacies that organizations encounter while implementing and managing these applications in such a dynamic setting.
RQ5
How to address security, automation, deployment, and monitoring in container-based applications within multi-cloud environment?
This question seeks to investigate solutions aimed at tackling the challenges of security, automation, deployment, and monitoring within container-based applications operating in multi-cloud environment. By providing actionable guidance and recommendations, the goal is to elevate the overall management and performance of these applications in the context of multi-cloud environment.
Tools and Frameworks
RQ6
What tools and frameworks are used to implement container-based multi-cloud applications?
To investigate and identify the specific tools and frameworks utilized in the implementation of container-based multi-cloud applications, providing an overview of the technological landscape and aiding developers and organizations in making informed decisions when selecting appropriate tools for their application deployment and management.
2.2. Study Search and Selection Process
The study search and selection process for this study is divided into two phases. In the first phase, we conducted a primary search using a specific search string on various databases (see Table 2 and the Summary sheet in (Waseem et al., 2024)). In the second phase, we applied the snowballing technique on primary studies that were selected from the first phase.
Table 2. Search string and targeted search area for databases
Search String
((multicloud OR multi-cloud OR multi cloud) AND
(container*) AND (automation) And (security) AND (monitoring) AND (deploy*))
Databases
Database Targeted search area
ACM Digital Library Paper title, abstract
IEEE Xplore Paper title, keywords, abstract
Scopus Paper title, keywords, abstract
SpringerLink Paper title, abstract
ScienceDirect Paper title, keywords, abstract
Wiley Online Library Paper title, abstract
Engineering Village Paper title, abstract
Web of Science Paper title, keywords, abstract
2.2.1. Primary Search
The primary search involved querying digital databases (see Table 2) using customized search string. The search strings were executed concurrently on eight databases, as illustrated in Figure 2, spanning between January 2013 and March 2023 when we started this SMS. We chose 2013 as the starting point because it is the year when container technologies like Docker emerged (Docker, 2023), marking a key phase in containerization and cloud computing. This period also saw a rise in relevant research to containers in multi-cloud environment, making it ideal to understand the development of containerized applications in multi-cloud environments over the past decade. We followed the steps mentioned below to select the relevant studies.
Table 3. Inclusion and exclusion criteria for selecting primary studies in this SMS
Selection Criteria Inclusion Criteria Exclusion Criteria
Language
English
Non-English
Study Type
Primary studies including peer-reviewed journal articles, book chapters, conference papers, workshops, and symposium papers
Secondary studies or non-peer-reviewed content (e.g., blogs, webpages, videos, white papers, technical reports, grey literature)
Study Focus
Studies addressing container roles and strategies, architectural patterns, quality attributes, adoption drivers, automation, deployment, monitoring, security challenges, solutions, and tools/frameworks for multi-cloud environment.
Studies that focus solely on traditional cloud computing, single-cloud environments, or unrelated topics
Study Duration
Studies published between January 2013 and March 2023
Studies published before January 2013 or after March 2023
  • Step 1: Extraction of Studies: We ran custom search strings in the selected databases (see Table 2) to retrieve study titles, author names, publication years, venues, publication types, and abstracts. This initial search yielded 5,676 studies from eight databases.Following initial retrieval, we did not apply Cohen’s Kappa as this step is automated and does not involve subjective decision-making.
  • Step 2: Title and Keyword Screening: This step involves reviewing the titles and keywords of the collected studies to assess their relevance to the research topic. Initially, we removed any duplicated studies obtained from different databases (such as IEEE Xplore and Scopus) by sorting them in ascending order. This resulted in the elimination of several hundred duplicate studies. We divided the remaining studies between two authors (R1, R2), who independently applied the inclusion and exclusion criteria to identify studies relevant to our research objectives, as detailed in Table 3. To assess the consistency of study selection and mitigate bias, a Cohen’s Kappa analysis (Cohen, 1968) was performed on a random sample of 40 studies, as illustrated in Formula 1:
    (1) $\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}$
    where $P_{o}$ represents the relative observed agreement among authors, and $P_{e}$ the hypothetical probability of chance agreement. The results, recorded in Table 4.a, indicated a Cohen’s Kappa of 0.531. This value reflects a moderate level of agreement between R1 and R2. Any disagreements were systematically resolved through discussions among all authors, leading to a consensus. Following this process, the number of studies was reduced to 809 for further analysis.
  • Step 3: Abstract-based Screening: During this step, we conducted a thorough review of the abstracts of the collected studies to determine their relevance to our research topic. Two researchers (R1 and R2) independently examined each abstract, assigning a status of ”relevant,” ”irrelevant,” or ”doubtful” with respect to our research goals. To reduce subjective bias and to quantify the consensus between the researchers, we conducted a Cohen’s Kappa analysis on a subset of 30 randomly selected studies. The outcomes of this assessment are depicted in Table 4.b, which show substantial agreement between R1 and R2, with 25 studies being unanimously considered relevant and only 4 studies disagree (3 with one researcher favoring relevance and the other not, and 1 where both disagreed) The Kappa value, calculated based on the observed agreements and disagreements was approximately K= 0.268 which suggests a fair agreement between the two researchers. Consequently, upon the compilation of step, we got 303 studies for more detailed scrutiny.
  • Step 4: Full-Text Screening: In this step, we reviewed the full texts of 303 studies that had been shortlisted through abstract-based screening. To ensure consistency in our review process, we randomly chose 20 studies (see Table 4.c) for a Cohen’s Kappa analysis to measure the agreement between two researchers. The resulting Kappa value was K=0.318, which represents a fair level of agreement. This measure confirms that our selection criteria were applied consistently during the full-text screening studies.
Table 4. Kappa Agreement across different screening steps.
(a) Title and Keyword Screening on 40 Randomly Selected Studies
R1 Yes R1 No
R2 Yes 35 2
R2 No 1 2
Total 36 4
(b) Abstract-based Screening on 30 Randomly Selected Studies
R1 Yes R1 No
R2 Yes 25 3
R2 No 1 1
Total 26 4
(c) Full-Text Screening on 20 Randomly Selected Studies
R1 Yes R1 No
R2 Yes 16 2
R2 No 1 1
Total 17 3
2.2.2. Snowballing
In Phase 2, we utilized the snowballing technique, as described in (Wohlin, 2014), to examine references within primary studies for identifying additional relevant studies. This strategy was augmented by forward snowballing, where we gathered studies that cited the selected studies, and backward snowballing, which involved using references within the selected studies. Notably, we encountered several dozen studies during the forward and backward snowballing that had been excluded in primary search. This phase resulted in the addition of 8 more studies, bringing the final count to 86.
2.3. Data Extraction and Analysis
Data Extraction: The data extraction form was designed based on predefined data items (see Table 5) that were formulated to address the RQs specified in Table 1. To ensure the reliability of the extracted data items, the pilot data extraction was conducted on ten studies by the first author, and all the other authors assessed the extracted data. Subsequently, the first author employed a revised set of data items (e.g., D11, D12, D16), determined after evaluating the extracted data items, for the formal data extraction from the selected studies. To mitigate the personal bias and ambiguity, all authors engaged in discussions regarding the extracted data. The data items labeled as D1 to D11 present a summary of the demographics of the primary studies selected, while D12 to D26 are specifically employed to address RQ1 to RQ6. A concise description of each data item is also presented in Table 5. Finally, Google Sheets were used to record and further analysis of the extracted data.
Data Analysis: We employed a descriptive statistics to analyze the quantitative data from data items D1, D5, D7, D8, and D9. For the remainder of the data, which primarily consist of qualitative free-text descriptions (e.g., study aim, roles of containers, challenges, and solutions), we conducted a thematic analysis in accordance with the guidelines outlined in (Braun and Clarke, 2006). Our thematic analysis process consists of the following steps:
  1. (1)
    Data Familiarization: We conducted a thorough review of the selected studies by repeatedly reading through them and meticulously noting key points related to study aims (D7), contributions (D8), roles of containers (D9), implementation strategies (D10), architectural patterns (D11), quality attributes (D12), tactics (D13), motivations (D14), and challenges and solutions (D15-D21) in automation, deployment, monitoring, and security challenges and solutions, as well as the tools, languages, and frameworks (D22) employed.
  2. (2)
    Generation of Initial Codes: After developing a thorough understanding of the data, we created an initial set of codes derived from the information extracted concerning the data items identified in the previous step.
  3. (3)
    Identification of Types and Emerging Themes: In this step, we conducted a two-tiered analysis. Initially, we examined the codes to ascertain their types. Subsequently, we developed subcategories based on these types, and then formulated overarching categories that encompass the related subcategories.
  4. (4)
    Critical Evaluation of Types, Subcategories, and Categories: All authors actively participated in rigorously reviewing and refining the coded data, including types, subcategories, and categories. During this collaborative process, we redefined, merged, or dropped certain themes based on collective input and discussion.
  5. (5)
    Defining and Naming Categories: At this point, we provided explicit definitions for each of the identified themes and further refined them, ensuring that the terminology used for the categories was precise and unambiguous.
Table 5. Data items to be extracted in this SMS
Code Data Item Description RQ
D1 Index The ID of the study Demographics
D2 Publication Year Publication year of the study
D3 Publisher The publisher of the study
D4 Venue The name of the publishing venue
D5 Publication Type Journal, conference, workshop, and book chapter
D6 Authors’ Affiliation The affiliation of the authors
D7 Study Aim The aim of the paper
D8 Study Contribution The contributions made by the paper
D9 Role of Containers Role of containers in the study RQ1
D10 Container Implementation strategy Strategy used for container implementation
D11 Architectural Pattern and Strategies Architectural patterns and strategies reported in the study RQ2
D12 Quality Attributes Quality attributes discussed in the study RQ3
D13 Tactics Tactics used in the study
D14 Automation Challenges Challenges in automation discussed in the study RQ4
D15 Automation Solutions Solutions for automation challenges discussed in the study RQ5
D16 Deployment Challenges Challenges in deployment discussed in the study RQ4
D17 Deployment Solutions Solutions for deployment challenges discussed in the study RQ5
D18 Monitoring Challenges Challenges in monitoring discussed in the study RQ4
D19 Monitoring Solutions Solutions for monitoring challenges discussed in the study RQ6
D20 Security Challenges Challenges in security discussed in the study RQ4
D21 Security Solutions Solutions for security challenges discussed in the study RQ5
D22 Tools, Languages, and Frameworks Tools, programming languages, and frameworks RQ6
The process of obtaining research themes executed in this SMS is shown in Figure 2. Two researchers participated in the process in order to reduce personal bias. The most important activity of this process is brainstorming sessions that were mainly conducted while reviewing, defining, and naming the research themes. In these sessions, both researchers discussed and validated the research themes found.
Concerning data item D9 (testing approaches) and D10 (testing challenges), we used the open coding and constant comparison techniques from Grounded Theory (Glaser and Strauss, 1967) to analyze the qualitative data extracted from the selected studies.
Finally, we provided a replication package (Waseem et al., 2024) containing the results of each phase of the study selection process (e.g., Phase 1, Phase 2) and detailed results (e.g., Contributions, Patterns, QAs, Challenges, Solutions) for verification and validation purposes of this SMS.
3. Results
This section presents the results of the SMS,derived from analyzing data from chosen studies. Section 3.1 provide a summary of the demographic data, classifications, and the mapping of research themes identified in the studies. Section 3.2 presents the results concerning container roles and implementation strategies. Section 3.3 reports on the patterns and strategies employed in containerized-based applications in multi-cloud environment. Section 3.4 presents the QAs and related tactics for containerized applications in multi-cloud environment. Section 3.5 presents the security challenges and solution framework, while Section 3.6 presents the automation challenges and solution framework. Section 3.7 presents the deployment challenges and solution framework, and Section 3.8 presents the monitoring challenges and solution framework. Lastly, the tools and frameworks are reported in Section 3.9.
3.1. Demographics, Research Aims, and Contributions
Studies Yearly Distribution: Figure 3-a illustrates the annual distribution of studies found in eight databases from January 2013 to March 2023. After applying our selection criteria, the final dataset includes studies published from 2014 to 2022. We found no primary studies on containerization in multi-cloud environments meeting our criteria in the first three months of 2023. The bars displays yearly study counts, with a trend showing the overall direction. A peak in 2017, followed by a decrease and occasional fluctuations, suggests an early spike in interest then more targeted research.
Figure 3. Demographics of the Selected Studies, Including: a) Frequency and Categories of Publications; b) Publisher of the Research; c) Affiliations of the Authors
Publishers Distribution: The pie chart presented in Figure 3-b illustrates the distribution of identified studies among different publishers. The largest portion of the studies, accounting for 72.09%, is published by IEEE, indicating its significant contribution to the field. Scopus follows with 11.62%, while Engineering Village, ACM, and Springer Link represent smaller portions at 6.97%, 5.81%, and 3.48%, respectively. The chart emphasizes the dominance of IEEE in the identified studies, suggesting its importance as a primary source of relevant literature.
Study Types: Figure 3-a provides a breakdown of the types of the selected studies. It is apparent that conferences dominate as the most common venue, comprising a substantial 72.09% of the total publications. Journals, which are often considered as venues for more mature and rigorously peer-reviewed research, make up 19.76%. This indicates a strong preference for disseminating information through conferences, which may be due to the faster publication process and opportunities for networking and collaboration. Symposiums and workshops account for 6.97% and 1.16% respectively, suggesting that these are less common avenues, possibly due to their more specialized or focused nature.
Authors’ Affiliations: Figure 3-c illustrates the distribution of authors’ affiliations in the collected publications. A staggering 93.02% of the affiliations are associated with academia. This overwhelmingly high percentage emphasizes the central role that academic institutions play in the production and dissemination of research in the field. Industry affiliations represent a mere 4.65%, which could imply that industry professionals may be more focused on practical applications and might not publish frequently in the venues considered. Mixed affiliations, at 2.32%, indicate a small proportion of collaborations between academia and industry. This low rate of cross-affiliation collaboration may suggest that there is potential for increased synergy between academic and industrial entities in future research endeavors.
Mapping of Research Themes: The thematic analysis of selected studies aims on containerization in multi-cloud environment, as illustrated in Figure 4, indicating various significant insights.The prominent research themes identified include “Scalability and High Availability” with 20 studies, and three other themes: ’Security and Privacy”m “Multi-Cloud Container Monitoring and Adaptation”, and “Performance and Optimization”, each attracting attention in 15 studies. This highlights that the core areas of interest in multi-cloud environment are proficient container management, scalability, and performance optimization. “Security and Privacy’ also stands as a considerable theme, as demonstrated by the 15 studies focusing on areas such as secure container handling, data security, and blockchain applications. This emphasizes the necessity of protecting data and applications in the context of containers spreading across multi-cloud environment.
Moreover, it is noted that several studies encapsulate more than one theme. For instance, study S19 is catalogued under both ‘Multi-cloud Orchestration and Management’ and ‘Application Deployment in Multi-cloud Container’, signifying that it encompasses aspects of both Kubernetes and microservices architecture.
In essence, the exploration into containerization in multi-cloud environment is substantially inclined towards, scalability, performance optimization security, and orchestration, with several studies showcasing inter-thematic relevance. This is suggestive of the diverse challenges and opportunities present in deploying containers within multi-cloud contexts.
Major Contributions of the Selected Studies: Table 6 provides an overview of the contributions made by the selected studies. The contributions are primarily categorized into five main categories and 22 subcategories (see the Contributions sheet in (Waseem et al., 2024)), through thematic analysis. The selected studies primarily focused on proposing and validating reference architectures, security approaches, cloud management and interoperability frameworks, orchestration and deployment approaches, as well as approaches and tools for optimization and deployment of containerized applications in multi-cloud environment.
The studies related to the reference architecture category mainly contributed to the understanding and design of architectures specifically tailored for multi-cloud containerization. They aimed to address the complexities associated with containerization in multi-cloud environment. On the other hand, the studies in the security approaches category discuss concerns related to cloud analysis, security and performance, cybersecurity, and access control within multi-cloud contexts.
Figure 4. Selected studies research aims in containerization for multi-cloud environment.
The studies on cloud management and interoperability frameworks primarily aim to improve resource and service management, monitoring, and deployment in multi-cloud environment. Their goal is to enhance the performance and efficiency of managing resources and services across multiple clouds. Meanwhile, the research in the area of orchestration and deployment approaches introduces methods for bench-marking, optimizing, and simulating container orchestration in multi-cloud settings. The objective of these studies is to streamline the process of deploying and orchestrating containers within multi-cloud environment.
Lastly, the research on optimization and deployment approaches and tools focuses on improving container deployment and creating efficient management solutions for multi-cloud environment. These studies underscore the importance of optimizing resource allocation and developing practical management tools for use in multi-cloud settings. Overall, these studies contribute to proposing reference architectures, security approaches, cloud management, and interoperability frameworks. These studies also offer strategies and tools for orchestration, deployment operations, and optimization in multi-cloud containerization.
Table 6. Selected Studies’ Contributions on Containerization in Multi-Cloud Environment
Category Subcategory Study #
Reference Architectures and Models for Cloud General Cloud Architectures S26, S33, S35, S42, S76
Fault Tolerance & Self-Healing S12, S55, S67, S74
Migration & Testing Architecture S17, S79
Domain-Specific Language S80, S82
Performance Modeling S28, S62
Optimization & Reliability Model S50, S66
IoT & Service Mesh Architecture S05, S59
Security Approaches Cloud Analysis Approach S18, S24, S37, S65, S69, S86
Security and Performance Solution S10, S23, S38, S75, S77
Cybersecurity & Access Control S11, S30, S34, S71, S73
Security Approach S06, S25, S40, S83
Cloud Management and Interoperability Framework Resource & Service Management Framework S48, S49, S47, S15, S58
Monitoring & Deployment Framework S19, S45, S46, S08
Interoperability Framework S03, S81
Evaluation & Modeling Framework S20, S85
Orchestration and Deployment Approach Benchmarking & Optimization Approach S01, S02, S41, S51, S68
Service Management Approach S07, S14, S29
Simulation & Integration Approach S27, S31, S56
Storage & Services Solution S64, S75
Optimization and Deployment Approaches and Tool Optimization & Scheduling Approach S22, S32, S36, S39
Multi-Cloud Platform S09, S16, S43, S60, S72
Deployment Approach and Tool S21, S52, S57, S61
[…]
3.2. Container Roles and Implementation Strategies (RQ1)
The next two subsections describe the roles and implementation strategiesS related to containers in multi-cloud environment.
3.2.1. Container Roles
Table 7 presents a systematic breakdown of the roles of containers in various aspects of multi-cloud environment. It is organized by the “Category Name” and the “Role of Containers” which detail the specific context or usage of containers in that category.
1. Container-based Application Deployment: This category discusses the use of containers for the deployment of various types of applications. This includes web apps, Platform as a Service (PaaS) solutions, data-intensive applications, and Internet of Things (IoT) apps, among others. Containers provide a means of packaging and distributing applications, including their dependencies, across various environments.
2. Container-based Service Hosting: This category refers to the use of containers as a hosting environment for different services. Container-based service hosting allow for scaling and isolation of these services, creating an efficient multi-cloud environment.
3. Container-Based Resource Management: This category discusses the use of containers to manage computational resources. Containers allow for better multi-tenancy solutions, resource sharing, and reduction in complexity, among other benefits.
4. Container Technology Features: This category discusses the specific features and advantages of container technology, such as lightweight virtualization, use in edge computing, and compatibility with orchestration tools like Kubernetes.
5. Containerization in App Development and IoT: The focus of this category is on the use of containers for app development, particularly in the context of IoT. Containers provide a means of encapsulating software units, increasing portability, improving efficiency, and integration with DevOps practices, among other benefits.
6. Performance and Efficiency through Containers: This category discusses how containers can enhance performance and efficiency in a cloud environment. Containers can improve scalability, isolation, and cost-effectiveness, and also aid in platform optimization.
7. Workload Management via Containers: This category is about the use of containers to manage computational workloads across multiple cloud environments. Containers facilitate the migration and execution of these workloads across different platforms.
8. Security and Compliance via Containers: In this category, containers are discussed as a tool for enhancing security and compliance in multi-cloud environment.
9. Other Container Uses: This category includes various other uses of containers, such as data storage, support in architecture, blockchain reliability, scheduling algorithm evaluation, and virtual network composition.
Each of these categories demonstrates a dedicated aspect of how containers can contribute to efficient and effective multi-cloud management. The studies listed in each row (e.g., S01, S02) provide a more in-depth discussion and evidence for the roles and applications of containers.
Table 7. Roles of containers in multi-cloud environment
Category Role of Containers ID
Container-based Application Deployment Containerized Web-App Deployment S01
Container-Driven App Deployment S05, S08
Containerized PaaS Prototyping S09
Data-Intensive App Deployment via Containers S28
Containerized App Management S29
Container-Based Aneka Service Deployment S51
Container-Driven Microservices Deployment & Scaling S62
Container-Based Multi-Cloud App Deployment & Scaling S74
Containerized IoT App Deployment S85
Container-based Service Hosting Container-Based App Component Hosting S02
Containerized Service Hosting S15
Container-Driven MicroCloud Scaling & Isolation S42
Container-Based Resource Management Container-Based Multi-Tenancy Solution S04
Container-Driven Resource Sharing & Multi-Tenancy S06
Multi-Tenant Service Management via Containers S17
Container-Enabled OS-Level Virtualization S40
Container-Based Complexity Reduction S41
Container Technology Features Container Usage in Edge Computing S22
Container-Enabled Lightweight Virtualization S24
Key Features of Container Technology S25
Container-Based Edge Computing Virtualization S35
Container Leverage in Kubernetes S70
Containerization in App Development & IoT App Development & IoT Integration via Containers S16
Container-Based OpenVAS & Integration of Chef S30
Containers as Standardized Software Units S32
Container Placement & Orchestration in Fog Computing S33
Container Deployment Speed & DevOps Integration S36
Containerizing Apps for Privacy & Compliance S37
Container-Driven Portability & Efficiency Improvement S39
Preventing Vendor Lock-In via Containers S63
Container-Based Extensible App Development S76
Performance & Efficiency via Containers Scalable App Building via Containers S18
Container-Driven SaaS Performance Isolation S23
Cloud Performance Enhancement via Containers S48
Evaluating Performance Impact of Containers S59
Cost Efficiency of Microservices via Containers S68
Container-Based Cloud Platform Optimization S69
Workload Management via Containers Container-Based Virtual Cluster Implementation S21
Multi-Cloud Workload Migration via Containers S60, S65
Seamless Workload Migration via Containers S65
Container-Based Multi-Cloud Platform Execution S79
Other Container Uses Container-Based Azure Data Storage S54
Container & VM Support in Architecture S55
Blockchain-Based Multi-Cloud Reliability via Containers S66
Evaluating Scheduling Algorithm via Containers S67
Container-Based Virtual Network Composition S72
[…]
3.2.2. Implementation Strategies
Table 8 provides an overview of various implementation strategies for container technology in multi-cloud environment. The strategies are grouped into several broad categories, each with a unique focus and varying subcategories. Within these subcategories, each implementation strategy is listed with a concise and revised name for reference.
1. Container Deployment and Management: This category focuses on how containers are deployed and managed in multi-cloud environment. An example from the “Deployment Strategies” subcategory is “Automated Orchestration for Web Applications”, where container deployment is automated for web applications. From the “Container Management” subcategory, the strategy “Container Image Optimization” highlights how the size and efficiency of a container can be optimized.
2. Container Orchestration: In this category, the focus is on coordinating and managing the life cycle of containers in a multi-cloud setup. An instance from the “Cloud Orchestration” subcategory is “SLA-Based Container Strategies”, illustrating how Service Level Agreements (SLAs) can guide the orchestration of containers. In the “Microservices Orchestration” subcategory, “Microservices Architecture in IoT Cloud Apps” represents a way to orchestrate microservices in IoT applications using containers.
3. Container Security: This category focus on the security of containers. An example from the “Container Placement” subcategory is “Secure Placement with Deep Learning”, which employs deep learning techniques to secure the placement of containers. From the “Security Policies” subcategory, the strategy “Access Control and Defense for Container Security” suggests ways to control access and defend containers from security threats.
4. Performance and Scaling: This category addresses the optimization of container performance and the scaling of container-based applications. For example, from the “Performance Optimization” subcategory, “LXC Virtual Cluster for Performance” shows how Linux Containers (LXC) can be used to enhance the performance of virtual clusters. An example from the “Scaling Strategies” subcategory is “Edge Data Streaming and Real-Time Database”, highlighting a strategy for scaling data streaming and real-time databases at the edge.
5. Cloud Service and Networking: This category deals with the management of cloud services and networking in containerized environments. “Multi-Tenant Service with Docker and Cloudant” from the “Cloud Service Management” subcategory illustrates how Docker and Cloudant can be utilized to manage multi-tenant services. From the “Networking” subcategory, the strategy “Edge Scheduling with Location-Awareness” proposes a way to schedule network services at the edge based on location.
6. Data Storage and Persistence: This category discusses strategies related to data management in containerized environments. For instance, the strategy “Firebase Cloud for Real-Time Database” presents an approach that uses Firebase Cloud to manage real-time databases in container-based applications.
[…]
Table 8. Container implementation strategies in multi-cloud environment
Category Subcategory Implementation Strategies ID
Container Deployment and Management Deployment Strategies Automated Orchestration for Web Applications S02
Kubernetes for Containerized Applications S05
Multi-cloud App Development Framework S07
Modular Approach to Container Implementation S08
Deployment with Docker Images S09
Docker Image Creation and Management S31
Container Management Container Image Optimization S32
Efficient Image Reorganization S36
Orchestrating with Docker and Kubernetes S37
Executable Images on Multi-cloud Platforms S79
Container Orchestration Cloud Orchestration SLA-Based Container Strategies S04
Task Division and Scaling in Container Strategy S12
Location-Aware Service Brokering S14
Docker Containers for Multi-tenant Services S17
Resource Acquisition for Aneka Containers S51
Cross-Cloud Container and VM Management S60
Container Orchestration Technologies Overview S33
Microservices Orchestration Microservices Architecture in IoT Cloud Apps S16
ADAPT Deployment of Microservices S27
Microservices Benchmark with Docker and Kubernetes S62
Container Security Container Placement Secure Placement with Deep Learning S06
Container Placement Strategies S39
Security Policies SLA-Oriented Performance Isolation S23
Access Control and Defense for Container Security S25
HIP Integration for Docker Security S77
Secure Connectivity in Multi-tenant Environment S72
Signed URLs for Access Control in Multi-cloud S71
Performance and Scaling Performance Optimization LXC Virtual Cluster for Performance S21
Scalable Multi-layer Architecture with Containers S28
Reactive Auto-scaling and Resource Provisioning S48
Istio for Control Plane Impact and Latency Reduction S59
Scaling Strategies Edge Data Streaming and Real-Time Database S35
MicroCloud Architecture for Efficient Scaling S42
Distributed Workloads with Kubernetes S70
Hybrid Abstractions for Container and Host S40
Cloud Service and Networking Cloud Service Management Multi-Tenant Service with Docker and Cloudant S17
Multi-Cloud Runtime Environment S76
Abstraction Layer for Cloud Service Standardization S63
Networking Edge Scheduling with Location-Awareness S22
Openstack Networking with Kuryr S55
Federated Framework for Cloud Storage S64
Host Identity Protocol for Docker Networking S77
Data Storage and Persistence Data Management Firebase Cloud for Real-Time Database S35
High Availability with SpyStorage MCSS S75
Data Storage with Object Store Service S76
3.3. Pattern and Strategies (RQ2)
Table 9 presents various patterns and strategies employed in containerized-based applications in multi-cloud environment. These patterns and strategies are divided into various categories and subcategories, each of which highlights different aspects of multi-cloud architecture, management, security, resilience, and migration patterns and strategies. Each category is briefly reported below.
1. Cloud Architectural Patterns and Models: This category reports different architecture patterns, communication and networking patterns, deployment patterns, and service models for containerized-based applications in multi-cloud environment. Architecture Patterns involve several design models for structuring the components of a system, providing a template that guides system designers in building complex systems. The leading patterns identified are microservice architecture and service-oriented architecture (SOA), both of which focus on building applications as suites of independently deployable services. Communication and Networking Patterns relate to how different services within the architecture communicate and network with each other. Service mesh and Service Chaining (SC) are the main patterns observed. Deployment Patterns pertain to different strategies for deploying software systems in multi-cloud environment. The two leading patterns are Blue-Green Deployment and Object Store Service. Service Models involve different ways of providing cloud services, with Platform-as-a-Service (PaaS) and Software-as-a-Servic (SaaS) leading the pack.
2. Cloud Management and Resource Allocation Strategies: This category presents the patterns and strategies related to the management of multi-cloud environment, container management, edge computing, IoT strategies, and resource management. For example, Multi-Cloud Management Strategies include various approaches for managing multiple cloud environments. Multi-Cloud Computing Strategy and Hybrid Cloud Architecture are found leading strategies in this category. Container Management Strategies involve the management of containers in multi-cloud environment, with the Linux Container (LXC) project and Container Engine as the primary patterns identified. Edge Computing and IoT Strategies pertain to the use of edge computing and IoT in multi-cloud environment. Mobile Edge Computing (MEC) and Fog Computing are leading strategies for container management in multi-cloud environment. Resource Management Strategies involve methods for managing resources in the multi-cloud with the help of Kubernetes and Minikube.
3. Cloud Security and Resilience Strategies: In this category, the emphasis is on strategies for ensuring the security and resilience of software systems in multi-cloud environment. Security and Resiliency Patterns are strategies to ensure the system is secure and can recover from failures. The two most prominent patterns are the security-by-design approach and encryption of files. Fault-tolerance Strategies are methods used to ensure that systems can continue to function even when there are failures. Fault-tolerance with Redundant Engines and Multi-cloud Systems Fault-Tolerant Workflow are the leading patterns.
4. Cloud Migration Strategies: This category involves patterns and strategies for moving applications and data between cloud environments and optimizing the migration process. Image-based Migration, for example, is an efficient way to migrate applications because it packages the applications along with their dependencies, configurations, and runtime environment. Migration to a virtualized container, such as Docker, is increasingly popular because it guarantees application portability across various environments and scalability on different cloud platforms.
Table 9. Patterns and strategies for container-based applications a multi-cloud environment
Categories Subcategories
Patterns
ID
Cloud Architectural Patterns and Models Architecture Patterns
Microservice Architecture
S02, S09, S16, S77, S80, S24, S31, S33, S68
Service-Oriented Architecture (SOA)
S51, S52, S86
Multitier Architecture, Layered Architecture
S02, S04, S54
Client-Server Architecture
S03, S54
Component-based Architecture
S22
Event-driven Architecture
S44
Multi-agent Architecture
S74, S48
Master-worker Nodes Architecture
S33
Netflix Zuul-based Seeker Component
S17
Multi-tenant Cloud Service Architecture
S17
Stateful Engine Architecture
S17
Communication and Networking Patterns
Service Mesh
S05
Service Chaining (SC)
S50
Multi-Cloud APIs
S71
Third-Party APIs
S71
Service Request Broker
S07
Publish and Subscribe Communication Protocol
S45
Synchronous Network Communication Protocols
S01
Asynchronous Network Communication Protocols
S01
Network Function Virtualization (NFV)
S50
Service Discovery
S76
Deployment Pattern
Blue-Green deployment
S27, S30
Object store service
S76
Multi-cloud Deployment Model
S10
Distributed Deployment Model
S10
Accelerated Particle Swarm Optimization for Deployment
S16
Service Models
Platform-as-a-Service (PaaS)
S24, S41
Software as a Service
S79
Infrastructure-as-a-Service (IaaS)
S24
Function-as-a-Service (FaaS)
S41
Cloud Management and Resource Allocation Strategies Multi-Cloud Management Strategies
Multi-cloud computing strategy
S78, S81
Hybrid cloud architecture
S78
Hybrid/Multi-cloud Approach
S05, S83
Introducing multi-cloud Middleware
S52, S53
MAPE-K control loop
S46, S48
Connecting to multiple cloud service providers
S76
Multi-cloud load balancing
S23, S76
Multi-Cloud BP provisioning
S84
Cross-level orchestration of cloud services
S84
Cross-level monitoring and adaptation of BPs
S84
Container Management Strategies
Linux Container (LXC) project
S24
Container engine
S24
Docker Container Images
S24
Lightweight virtualization
S24
Portable application packaging
S24
One-container-per-app approach
S24
Data volumes and data volume containers
S24
Edge Computing and IoT Strategies
Mobile Edge Computing (MEC)
S22, S70, S43, S54
Fog computing
S32
Edge Services
S35
Connecting IoT Edge Devices to the Cluster
S05
Cloud of Things (CoT)
S05
Resource Management Strategies
Resource allocation based on different SLA levels
S23
Container Orchestration, Kubernetes, Minikube
S24, S29, S33
Cloud Security and Resilience Strategies Security and Resiliency Patterns
Security-by-Design approach
S51
Encryption of files
S71
Standardized APIs for file transfer
S71
Centralized/External ACM Service
S71
Distributed ACM Service
S71
Dynamic Switching between Authentication Methods
S71
ACM based on Signed URLs
S71
Attribute-Based Encryption and Signature
S75
Secure data sharing
S75
Data access control
S75
Local encryption and signing
S75
Authorized Tokens
S75
Byzantine quorum protocol
S75
Fault-tolerance Strategies
Fault-tolerance with Redundant Engines
S17
Multi-cloud systems fault-tolerant workflow
S67
RAFT Consensus Algorithm
S13
Cloud Migration Strategies
Service-oriented Migration
S15
Application-centric Migration
S15
Image-based Migration
S15
Migration to a Virtualized Container
S15
[…]
3.4. Quality Attributes and Tactics (RQ3)
Table 10 provides a comprehensive overview of the quality attributes and related tactics for containerized applications in multi-cloud environment. It is based on data extracted from 86 selected studies, which are organized following the ISO 25010 standard. This standard defines high-level QAs along with their sub-characteristics. Table 10 lists nine QAs, namely Performance (Efficiency), Security, Compatibility, Scalability, Reliability, Portability, Flexibility, Maintainability, and Usability. Each QA has its related terms, associated study numbers, and corresponding tactics to achieve them.
In Table 10, five tactics are provided for each quality attribute, due to the limited space. The complete set of 195 identified tactics is available in the provided dataset (Waseem et al., 2024), presenting a comprehensive map of tactics with QAs. However, it is important to note that these tactics are scattered across different studies, and only a few studies explicitly mentioned these tactics for specific QA.
Below we briefly discuss how each QA can be achieved by applying the mentioned tactics.
Performance (Efficiency): The quality attribute of performance (or efficiency) in containerized applications can be achieved by employing tactics such as efficient resource optimization and allocation, use of ML techniques for performance optimization, utilization of container technology, location-aware service brokering, and performance-oriented Service Level Agreements (SLA). For instance, optimizing resources ensures that the system uses the minimum possible resources while delivering the required output. Using ML techniques for performance optimization can enhance system performance by learning and adjusting the operational parameters. Similarly, utilizing container technology can help maintain optimal performance by encapsulating the application and its environment.
Security: Security can be enhanced using tactics like encryption and strong authentication mechanisms, implementation of firewalls, VPNs, or SDNs for network security, utilization of ML techniques to detect and prevent security threats and attacks, deployment of applications on diverse cloud providers, and consideration of users’ security specifications. For example, encryption provides a secure way of transmitting data, and ML can detect unusual behavior that could signal a threat or attack.
Compatibility: The tactics that can be employed to achieve compatibility include implementation of lightweight communication protocols and modes, standardization of interfaces for seamless integration, componentization and modular design for interoperability, prototyping and exploration of interoperability approaches for multi-cloud deployment, and interoperability standardization and federation between clouds. Standardizing interfaces can ensure that different software components can interact with each other seamlessly, enhancing compatibility.
Scalability: Scalability can be achieved through tactics like combining containerization and microservices for enhanced scalability, implementation of elastic resource allocation, enabling migration between multi-cloud services, combining container-based deployment with runtime monitoring and optimization, and leveraging orchestration, federated networks, and geographic placement. Elastic resource allocation ensures that the system can seamlessly scale up or down in response to changing demands.
Reliability: Tactics such as deploying redundant engines for fault tolerance, distributing resources and replicating applications for improved response time, enforcing redundancy and distributed services for availability, deploying parallel search and multi-cloud distribution, and building fault-tolerant systems can enhance system reliability.
Portability: Portability can be achieved using tactics such as virtual machine-based packaging, Docker image-based packaging, cloud-portable containerization, dynamic cross-level adaptation and provisioning, and adaptive rule-based system modification. These tactics ensure that an application can be easily transferred from one computing environment to another.
Flexibility: Flexibility can be improved through tactics like service-oriented architecture, on-demand dynamic allocation of resources across different cloud platforms, containers, container orchestration, model-driven development, risk analysis, and cross-cloud service orchestration. Using multi-cloud environment, for instance, offers the flexibility to choose services from different providers as per specific needs.
Maintainability: Tactics like microservices for easier maintenance, Infrastructure Provisioning as Code (IaC), containerization and image management, configuration management and templating, version control and change management, and continuous integration and deployment automation can enhance maintainability. For example, IaC allows developers to manage infrastructure more efficiently and minimize human errors, thus increasing maintainability.
Usability: Usability can be enhanced by employing tactics like a feedback-loop controller for multi-cloud infrastructure, container orchestration (e.g., Kubernetes), user-centric interface design, responsive and adaptive user experience, accessible design and compliance, and error handling and feedback mechanisms. For instance, a feedback-loop controller is effective in a multi-cloud context as it dynamically adjusts resources across different cloud platforms based on user interaction, ensuring a seamless and responsive user experience tailored to the multi-cloud setup.
[…]
Table 10. Quality attributes and tactics for containerized applications in multi-cloud environment
Quality Attribute Related Terms ID Tactics
Performance (Efficiency) Time behavior, Resource utilization, Capacity, Throughput, Response time S01, S02, S03, S06, S09, S14, S15, S16, S20, S21, S23, S25, S27, S29, S34, S35, S36, S37, S38, S39, S41, S43, S44, S47, S49, S52, S54, S57, S58, S59, S60, S62, S64, S65, S66 Efficient resource optimization and allocation
Utilizing machine learning techniques for performance optimization
Container technology utilization
Location-aware service brokering
Performance-oriented Service Level Agreements (SLA)
Security Confidentiality, Authenticity, Access control, Authorization, Privacy, Trustworthiness, Integrity S01,S05, S06, S09, S15, S23, S24, S25, S29, S34, S37, S38, S40, S44, S47, S52, S54, S55, S57, S59, S61, S64, S65, S66, S69, S71, S72, S73, S77, S86 Encryption and strong authentication mechanisms
Implementation of firewalls, VPNs, or SDNs for network security
Utilizing machine learning techniques to detect and
prevent security threats and attacks
Deployment of cloud applications on diverse cloud providers
Consideration of users’ security specifications and
addressing CSP incompatibilities
Compatibility Interoperability, Co-existence S15, S16, S21, S24, S27, S42,S44, S45, S49, S63, S64, S71, S77, S80, S82 Implementing lightweight communication protocols and modes
Standardization of interfaces for seamless integration
Componentization and modular design for interoperability
Prototyping and exploring interoperability approaches for
multi-cloud deployment
Interoperability standardization and federation between clouds
Scalability Capacity, Extensibility, Elasticity, Throughput, Responsiveness S01, S08, S09, S14, S15, S16, S17, S18,S19, S21, S22, S24, S25, S27, S29, S31, S35, S41, S42, S44, S45, S54, S57, S62, S66, S69, S72, S74, S76, S77, S82, S84
Combining Containerization and Microservices for enhanced scalability
Implementing Elastic Resource Allocation
Enabling Migration between Multi-Cloud Services
Combining Container-based Deployment with Run-time Monitoring
and Optimization
Leveraging Orchestration, Federated Network,and Geographic Placement
Reliability Maturity, Availability, Fault tolerance, Recoverability S01, S02, S04, S10, S13, S17,S19, S20, S22, S25, S27, S32, S33, S35, S37, S44, S47, S51,S53, S54, S55, S56, S57, S60, S64, S65, S66, S67, S71, S73, S74, S75, S82 Deploying Redundant Engines for Fault-Tolerance
Distributing Resources and Replicating Applications for Improved
Response Time
Enforcing Redundancy and Distributed Service for Availability
Deploying Parallel Search and Multi-Cloud Distribution
Building Fault-Tolerant Systems
Portability Installability, Replaceability, Adaptability S18, S24, S25, S31, S62, S77 Virtual Machine-Based Packaging
Docker Image-Based Packaging
Cloud-Portable Containerization
Dynamic Cross-Level Adaptation and Provisioning
Adaptive Rule-Based System Modification
Flexibility Functional completeness, Functional correctness, Functional appropriateness S16, S21, S27, S31, S41, S61, S60, S70, S76, S84 Service-oriented Architecture
Dynamic allocation of resources across different cloud platforms based on demand
Container orchestration
Model-Driven Development, Risk Analysis
Cross-cloud service orchestration,
Maintainability Modularity, Reusability, Analyzability, Modfiabality, Testability S08, S17, S25, S27, S62, S74, S82 Microservices, easier maintenance
Infrastructure Provisioning as Code (IaC)
Containerization and Image Management
Configuration Management and Templating
Version Control and Change Management
Continuous Integration and Deployment Automation
Usability Appropriateness recognizability, Learnability, Operability , User interface aesthetics, Accessibility S09, S17, S27, S27, S37, S74 Feedback-loop controller for multi-cloud Infrastructure
Container Orchestration (e.g., Kubernetes)
User-Centric Interface Design
Responsive and Adaptive User Experience
Accessible Design and Compliance
Error Handling and Feedback Mechanisms
3.5. Security Challenges and Solution Framework (RQ4 and RQ5)
In an era of rapid technological advancement, ensuring data security is of paramount importance. In this section, we propose a framework designed to address the security challenges in containerized and multi-cloud environment. This proposed framework, composed of seven categories of security challenges and solutions illustrated in Table 11, is derived through a comprehensive review of existing literature. It covers a broad spectrum of categories, each capturing specific security challenges and their corresponding solutions. Each category is briefly described below:
Data Security in container-based applications category covers a wide range of challenges and solutions, including data protection, database security, data compliance, and secure data transfer. Our identified solutions revolve around the use of encryption and secure container orchestration, complying with regulatory requirements such as GDPR, HIPAA, and secure data management strategies. Furthermore, identified solutions emphasizes the role of security measures such as access control and secure transmission protocols in ensuring data privacy, service continuity, fault tolerance, and the secure transfer of data between different cloud environments.
Access and Communication Control category emphasizes maintaining secure multi-cloud connectivity and managing user access control. The identified solutions involve implementing encryption and authentication mechanisms, using secure communication protocols, and executing operations in a provider-agnostic manner to maintain security.
Container Security category addresses issues like co-resident security, container isolation, and providing host security. Implementing secure placement strategies for containers across multi-cloud platforms, utilizing container isolation techniques, and enhancing manual defenses to mitigate risks of compromised containers are some of the identified solutions in this category.
Infrastructure and Deployment Management Security category emphasizes the significance of infrastructure and deployment management security. It encompasses aspects such as secure data deployment, infrastructure orchestration, storage security, and migration security. Essential solutions encompass the application of encryption techniques, implementation of access control mechanisms, adoption of standardized and secure infrastructure orchestration practices, and ensuring the secure migration of containerized applications across multi-cloud environment.
Security Monitoring and Breach Prevention category addresses a range of challenges and solutions related to enhancing security. This includes the implementation of automated security assessments, the use of vulnerability scanners, and the establishment of automated bootstrapping mechanisms for Virtual Machines (VMs) that incorporate security configurations and policies. Additionally, it involves deploying Security Information and Event Management (SIEM) systems, as well as Intrusion Detection System/Intrusion Prevention System (IDS/IPS) tools. The category also covers the application of log analysis solutions to proactively detect and prevent security breaches.
Trust and Compliance Management category identified solutions encompass on implementation of robust security controls, adherence to industry best practices and standards, and establishment of trust-based coalitions to foster secure community formation.
Placement Strategy: It recommends devising a security model for services hosted on volunteering fogs within multi-cloud environment, including formulation of secure communication protocols and implementation of access control mechanisms.
[…]
Table 11. Security challenge and solution framework for containerized applications in multi-cloud environment
ID Challenge Challenge Description Solution Solution Description
Category 1: Data Security
S17 Data Protection Protecting data against threats in containerized multi-cloud setups. Multi-Cloud Data Protection through Secure Containers Deploys sensitive data in encrypted, isolated containers, using secure orchestration tools like Docker and Kubernetes.
S29 Database Security Securing databases within containerized multi-cloud environments. Enhanced multi-cloud Database Security via Containerization Secures databases by implementing encryption (e.g., Transparent Data Encryption) and robust access controls (e.g., RBAC, ABAC).
S37 Data Compliance Ensuring data processing and storage compliance across multi-cloud platforms. Multi-cloud Legal Compliance and Data Protection Maintains privacy and service quality by adhering to regulations (e.g., GDPR, HIPAA) and employing data protection strategies like encryption and audits.
S38 Data Security Adhering to data security standards in multi-cloud containerized systems. Secure CSP Interoperability and Selection in multi-cloud Tackles CSP compatibility issues and enhances security by implementing data handling practices and selecting compliant CSPs.
S53 Data Ownership and Privacy Tackles data ownership, privacy, and availability issues in multi-cloud setups. Multi-cloud Security via NoMISHAP Service Abstraction Uses NoMISHAP for service abstraction, enhancing privacy with encryption and pseudonymization, and ensuring availability with fault tolerance strategies like backups and failovers.
S58 Data Transfer Security Secures data during transfers in multi-cloud environments. Secure multi-cloud Integration via OpenStack Standardization Standardizes OpenStack APIs to prevent vendor lock-in, ensuring safe data transfers with encryption and secure protocols.
S85 Cloud Data Management Manages cloud data security, privacy, and compliance. Multi-cloud Data Security via DRA Framework Implements the DRA framework for comprehensive data management, ensuring encryption, access control, and compliance with standards.
Category 2: Access and Communication Control
S05 Connectivity and Access Control Ensures secure multi-cloud connectivity and robust user access control. Multi-cloud Containerized Application Security Implements TLS, JWT, OAuth 2.0 for secure authentication and authorization in multi-cloud setups.
S43 Secure Communication Secures communications across multi-cloud and MEC platforms. Middleware Secure Deployment for multi-cloud Uses middleware for secure cloud app deployment, with SSL/TLS and access control for secure communication.
S51 Provider-Agnostic Operations Facilitates secure, provider-agnostic operations in multi-cloud environments. Aneka-Based Security and Performance for Containerized multi-cloud Applications Employs Aneka for security and performance, offering transparent access and dynamic provider selection, while ensuring security via secure channels and monitoring.
Category 3: Container Security
S06 Co-resident Security Addresses security risks from container virtualization in multi-cloud environments. Secure Placement Strategy via Deep RL Uses deep reinforcement learning, like Kubernetes Cluster Autoscaler with RL, for secure container placement across multi-cloud platforms.
S25 Container Isolation Aims to prevent information leaks and power attacks through effective container isolation. Power and Leakage Management for multi-cloud Security Manages security via container isolation (e.g., Docker isolation) and Linux kernel-level resource isolation (e.g., SELinux).
S40 Container and Host Security Enhances security against compromised containers and strengthens defenses. Host-Container Cooperation for multi-cloud Defense Develops container-host defense strategies, utilizing tools like Falco, Sysdig, and practices for host hardening.
Category 4: Infrastructure and Deployment Management Security
S20 Data Deployment Secures private data deployment in multi-cloud environments. Secure Private Data Deployment Strategy Utilizes encryption, RBAC, and secure protocols (e.g., AES-256, SFTP, HTTPS) for safeguarding data deployment.
S44 Infrastructure Orchestration Orchestrates secure infrastructure for containerized applications. Standard Interface Implementation for Serverless Security Uses standard interfaces (e.g., OpenStack APIs) for secure, interoperable serverless application deployment across clouds.
S47 Storage Security Enhances security for multi-cloud storage in containerized setups. Multi-cloud Storage Security with MSSF Employs the MSSF for secure storage management, implementing RBAC, data encryption, and regular audits.
S80 Migration Security Ensures the secure migration of applications in multi-cloud environments. Infrastructure-Aware and Agnostic Secure Migration Leverages strategies and tools for safe migration, including image signing and data encryption, adhering to best practices like least privilege.
Category 5: Security Monitoring and Breach Prevention
S30 System Vulnerabilities Addresses vulnerabilities and patching delays in multi-cloud environments. Automated Security Evaluation and Bootstrapping Utilizes tools like OpenVAS and Nessus for security assessments, and automates VM setup with Terraform, Chef, Ansible, or Puppet.
S45 Breach Detection and Prevention Boosts breach detection and prevention in multi-cloud infrastructures. Metric Filtering and Fault Recovery for multi-cloud Security Enhances security with metric filtering, fault recovery, SIEM, IDS/IPS, and log analysis for proactive breach management.
Category 6: Trust and Compliance Management
S52 Application Security and Compliance Ensures security and compliance for applications across multi-clouds. Risk Analysis and Security SLAs with MUSA Framework Uses MUSA for risk analysis and security SLA management, incorporating secure coding, vulnerability management, and compliance audits.
S49 Data Security and Standardization Standardizes data security and cloud interfaces in multi-cloud environments. OpenStack API-based multi-cloud Security and Interoperability Utilizes OpenStack API for secure, standardized cloud interoperability, applying encryption, access control, and CIS benchmarks.
S57 Network Security Secures networks and trust in multi-cloud environments. Inter-cloud Communication Protection with SFC Employs Service Function Chaining (SFC) for secure inter-cloud communications, using encryption, integrity checks, and protocols like IPSec, WireGuard.
S83 Trust and Community Formation Builds secure communities and trust in multi-cloud settings. Trust-based Hedonic Coalitions for multi-cloud Security Forms trust-based coalitions for community security, leveraging IAM and trust frameworks for collaboration in multi-clouds.
S84 Orchestration and Adaptation Manages orchestration and adaptation for security and performance. Multi-cloud Service Orchestration and Cross-Level Adaptation Implements multi-cloud orchestration with tools like Kubernetes, Istio, and Prometheus for secure, adaptive application management.
Category 7: Placement Strategy
S33 Placement Strategy Secures service deployment on volunteer fogs in multi-cloud setups. Security Model for Volunteering Fog Services Creates a security model for fog-deployed services, incorporating SSL/TLS for communication and certificate or token-based access control.
3.6. Automation Challenges and Solution Framework (RQ4 and RQ5)
Table 12 presents an overview of the automation challenges and solution framework for containerized applications in multi-cloud environment. The figure consists of eight categories, each focusing on different aspects of automation challenges and their solutions in a multi-cloud environment.
Multi-Cloud Automation: This category reports challenges and solutions related to automating processes for or combining and administering various cloud computing environments. It addresses challenges such as multi-cloud integration and provisioning, and provides solutions such as proxies for environment interface and model-driven engineering techniques.
Automation in Deployment and Scaling: This category reports challenges and solutions related to automating the deployment and scaling processes of containerized applications across multi-cloud platforms. It tackles challenges like container placement automation and application runtime switch automation. Solutions include intelligent container placement algorithms and service drivers for automated runtime switch.
Resource Management Automation: This category reports challenges and solutions related to automating the management of resources in multi-cloud environment. It addresses challenges such as resource overbooking management and containerized applications management. Solutions include ML-based resource overbooking detection and custom Kubernetes schedulers.
Data and Application Migration Automation: This category reports challenges and solutions related to automating the migration of data and applications in multi-cloud environment. It covers challenges such as data integrity during migration and legacy web applications migration. Solutions encompass the implementation of self-sufficient management cycles for streamlined deployment, alongside the utilization of adaptable architectural patterns designed for the modernization of legacy applications within a containerized multi-cloud environment.
Testing and Benchmarking Automation: This category reports challenges and solutions related to the automation of testing and benchmarking processes for containerized applications in multi-cloud environment. It addresses challenges like user-oriented testing automation and benchmarking automation. Solutions include container-based lightweight virtualization for testing and smart Docker benchmarking orchestrators.
Standardization and Interoperability Challenges: This category reports challenges and solutions related to standardization and interoperability in the automation of containerized applications across multiple cloud environments. It covers challenges like standardization in automation processes and cloud provider selection. Solutions include service mesh networks and AI-driven cloud provider selection tools.
Application and Service Management: This category reports challenges and solutions related to the automation of application and service management processes in containerized multi-cloud environment. It addresses challenges such as 5G ecosystem component selection and end-to-end tail latency management. Solutions include open-source software components for customizable orchestration and modeling approaches for performance evaluation.
Runtime and Service Discovery: This category deals with challenges and solutions related to runtime management and service discovery in multi-cloud environment, which involves managing the complexity of network control and administration within containerized applications across multiple cloud environments.
Table 12. Automation challenge and solution framework for containerized applications in multi-cloud environment
ID Challenge Challenge Description Solution Solution Description
Category 1: Multi-Cloud Automation
S10 Multi-Cloud Integration Automating processes to manage the complexities of integrating multiple cloud environments Proxies for Environment Interface Proxies for individual environment interface, unified view of resources, shielding from underlying complexity in multi-cloud environment
S20 Multi-Cloud Provisioning and Adaptation Automating provisioning, deployment, and dynamic adaptation of multi-cloud applications for efficiency and timely response Model-driven Engineering Techniques Model-driven engineering techniques for dynamic provisioning, deployment, and adaptation in multi-cloud environment
S33 Fog Creation in Multi-Cloud Automating the dynamic creation of nearby fog in multi-cloud environment Dynamic On-demand Fog Computing Dynamic on-demand fog computing architecture in multi-cloud environment
S47 Intelligent User-Friendly Automation Addressing the lack of intelligent, user-friendly automation in delivery platforms MSSF MSSF (Multi-cloud Storage Selection Framework) for automatic selection of multi-cloud redundancy storage based on user needs and security rules
S44 Multi-Stack Orchestration Automating the specification, coordination, and orchestration of different infrastructure stacks and cloud services TOSCA-based Serverless Deployment Topology and Orchestration Specification for Cloud Applications (TOSCA) for automated deployment of multi-cloud serverless applications
S34 Secure Multi-Cloud App Creation Providing tools and mechanisms for non-security expert developers to automate secure multi-cloud app creation MUSA Framework MUSA framework for application modeling, risk analysis, cloud service selection, composite SLA generation in multi-cloud environment
S37 Unified Abstraction for Multi-Cloud Deployment Implementing a common abstraction for products offered in automated multi-cloud deployments Application-level Resource Managers Application-level resource managers, native multi-cloud support for enterprise applications in multi-cloud environment
S52 Multi-Cloud Deployment Integration Addressing integration complexity in multi-cloud deployment automation MUSA Deployer Tool MUSA Deployer tool for automation of deployment process, deployment plan, and security compliance in multi-cloud environment
S57 Deployment Across Federated Clouds Automating deployment across federated cloud providers Orchestrated Deployment Subsystem Orchestrated deployment subsystem, Beacon Broker, and federated environments for coordinated application deployment in multi-cloud environment
S58 Hybrid/Multi-Cloud Setup Automating the setup of hybrid/multi-cloud environment to address security, vendor lock-in, and interoperability issues Interoperability Standardization Interoperability standardization and federation of clouds in multi-cloud environment
S80 Multi-cloud Interoperability Facilitating multi-cloud interoperability, application topology definition/composition, and elastic runtime adaptation for containerized services Multi-cloud Interoperability Multi-cloud interoperability, application topology definition/composition, and elastic runtime adaption in multi-cloud environment
S81 Multi-cloud Brokerage System Overcoming API heterogeneity and promoting semantic interoperability for a brokerage system in multi-cloud container management. Managing Large Number of Cloud Services Interoperability and vendor lock-in caused by the diversity of cloud solutions, complexity of multi-cloud systems administration
S84 Multi-cloud Orchestration Implementing multi-cloud business process provisioning and cross-level orchestration for cloud services in a containerized setting Manual Migration to the Cloud Manual migration of business processes to the Cloud, limited multi-cloud and cross-cloud orchestration
S85 DevOps for Multi-cloud Building a DevOps reference architecture with a CI broker for automation within a containerized multi-cloud environment Agile Software Development Automation Automation of agile software development, application deployment using CI-broker, automated faster application delivery, automated and integrated testing of the application in multi-cloud environment
Category 2: Automation in Deployment and Scaling
S27 Multi-Cloud Deployment and Monitoring Tackling challenges in automating deployment and monitoring of applications across multi-cloud platforms Multi-cloud Strategy Multi-cloud strategy for risk distribution and provider dependency reduction in multi-cloud environment
S06 Container Placement Automation Addressing the complexity of automating container placement in multi-cloud environment Intelligent Container Placement Develop a machine learning-based optimization algorithm for intelligent container placement across multiple cloud environments, focusing on load balancing and minimizing network latency.
S07 Application Runtime Switch Automation Enabling automated application runtime switch among various cloud services Service Driver Service driver, multi-cloud application service access model, runtime service update, development framework in multi-cloud environment
S15 Fine-Grained Automation at Component Level Implementing autonomic fine-grained solutions at the service component level for better automation Mobility Service and Replication Mobility service, migration and replication of critical software components, autonomic fine-grained service migration/replication in multi-cloud environment
S28 DevOps Automation Ensuring effective automation of software development and operations management DevOps Development Model DevOps development model, continuous development, automated tools and management systems, autonomic management system, performance models, systematic decision-making in multi-cloud environment
S29 Service Deployment Automation Addressing manual deployment issues by automating service deployment Automation Scripts Automation scripts, automated deployment, and error reduction in multi-cloud environment
S30 Deployment Time Security Automation Implementing security automation during deployment time Automated Security Measures Automated security measures, security evaluation and bootstrap, vulnerability scanning, firewall deployment, Chef integration in multi-cloud environment
S31 Distributed Simulation Automation Automating distributed simulation environments in cloud infrastructures DevOps Methodologies DevOps methodologies, automated developer operations, rapid microservices deployment, containerization paradigm, DEVS Simulation Farm, Docker technology, SOA Microservices paradigm, DEVSML Stack in multi-cloud environment
S32 Software Deployment Speed Improvement Addressing slow software deployment issues by implementing effective automation, enhancing Quality of Experience (QoE) Reorganized Container Images Reorganized container images, modified Docker deployment process, asynchronous image layer retrieval, profiled container execution, overlay file system in multi-cloud environment
S36 Container Startup Automation Identifying essential files for container startup automation Profiled Container Execution Profiled container execution, essential file determination, special image layer creation in multi-cloud environment
S38 Dynamic Scaling of Resources Automating frequent deployment and dynamic scaling of resources in multi-cloud environment for improved efficiency and performance CSP Automation Capability CSP automation capability for user agility in containerization of multi-cloud environment
S24 Container Deployment and Scaling Implementing automated container deployment and scaling based on resource demand across multiple cloud platforms Kubernetes Kubernetes for automated container deployment and multi-cloud scaling in multi-cloud environment
S48 Elastic Resource Provisioning Implementing elastic resource provisioning for fluctuating business workloads in a containerized multi-cloud environment through automation Elastic Resource Provisioning Design a predictive analytics system coupled with automation tools for real-time, elastic resource provisioning, addressing fluctuating business workloads in a containerized multi-cloud environment
S50 VF Placement and Service Deployment Automating optimal Virtual Function (VF) placement and dynamic service deployment in multi-cloud environment VF Placement Optimization Optimization model for virtual function (VF) placement and dynamic VF allocation in multi-cloud environment
S68 Container and VM Deployment Orchestrating and automating the deployment of containers and VMs in multi-cloud environment Multi-Cloud Orchestration Develop a comprehensive multi-cloud management platform for streamlining orchestration and automation of containers and virtual machine deployment across different cloud environments
Category 3: Resource Management Automation
S04 Resource Overbooking Management Automating the management of resource overbooking in service containers across multi-cloud platforms ML-based Resource Overbooking Detection Machine learning-based resource overbooking detection in multi-cloud environment
S21 VCluster Management Automating cluster creation, deployment, and management for multiple Virtual Clusters (Vclusters) across diverse application frameworks ClaaS Service Model ClaaS service model, virtualized cluster environment, easy deployment and management in multi-cloud environment
S39 Containerized Applications Management Automating deployment, scheduling, and management of containerized applications in multi-user and multi-cloud environment Custom Kubernetes Scheduler Custom Kubernetes scheduler (label-affinity-scheduler) for container orchestration in multi-cloud environment
S41 VM Types Selection in Data Centers Addressing the large and complex search space of different VM types in various data centers through automation GA-based Algorithm GA-based algorithm, domain-tailored service clustering method for complexity reduction in multi-cloud environment
S42 Resource Commissioning and De-Commissioning Speeding up resource commissioning and de-commissioning through automation for timely system scalability MicroCloud Architecture MicroCloud architecture, fine-grained resource allocation, TOSCA library, meta-workflow, coordinated adaptation workflow in multi-cloud environment
S43 Cloud Federation Automation Implementing automation support for federation or inter-platform portability in multi-cloud environment Middleware Platform Middleware platform for security, cost, and performance optimization, data locality, privacy needs in multi-cloud and MEC environments
S45 Resource Provisioning and Elasticity Detection Automating resource provisioning for cloud services, monitoring runtime configurations, and detecting elasticity actions JCatascopia JCatascopia, a fully-automated and modular cloud monitoring framework with elasticity actions, metric rule mechanism, and low-level monitoring metrics in multi-cloud environment
S51 Resource Acquisition Automation Automating resource acquisition processes for applications to function without human intervention Aneka Platform Aneka platform for automatic acquisition and release of resources from multiple clouds in multi-cloud environment
S74 SLA Enforcement Automation Automating the testing and corrective actions for SLA enforcement in containerized multi-cloud environment to reduce expense and time consumption SLA Compliance Testing Manual testing and corrective actions for SLA compliance are expensive and time-consuming in multi-cloud environment
Category 4: Data and Application Migration Automation
S01 Data Integrity during Migration Safeguarding data integrity during the automation of migration processes in multi-cloud environment Autonomous Management Cycle Autonomous management cycle for efficient deployment in multi-cloud environment
S17 Legacy Web Applications Migration Facilitating the automated migration of legacy web applications to cloud services Reusable Architectural Pattern Reusable architectural pattern for containerization in multi-cloud environment
S69 Application Porting to Cloud Automating the porting of existing applications to cloud infrastructures CloudSME Simulation Platform CloudSME Simulation Platform for fast development and deployment of complex multi-cloud applications, commercial accounting, and billing
Category 5: Testing and Banchmakring Automation
S09 User-Oriented Testing Automation Accelerating user-oriented software testing in diverse and heterogeneous environments through automation S Container-based Lightweight Virtualization Container-based lightweight virtualization, just-in-time deployment, streamlined multicomponent prototyping in multi-cloud environment
S02 Benchmarking Automation Implementing an automated benchmarking process for cloud hosts from diverse providers in multi-cloud environment Smart Docker Benchmarking Orchestrator Smart Docker Benchmarking Orchestrator for containerization in multi-cloud environment
S13 CI for Dependable Automation Leveraging Computational Intelligence (CI) to enhance dependability in automated multi-cloud containerized environments Dependability Enhancement with CI Leverage Computational Intelligence (CI) to enhance dependability in automated multi-cloud containerized environments
Category 6: Standardization and Interoperability Challenges
S05 Standardization in Automation Implementing standardization in the automation processes for containerized applications across multiple cloud environments Service Mesh and Multi-cloud Federation Service mesh network and multi-cloud cluster federation for containerization in multi-cloud environment
S70 Cloud Provider Selection Developing optimal strategies for cloud provider and node configuration selection to balance cost and runtime in a multi-cloud, containerized environment Optimal Cloud Provider Selection Create an AI-driven decision-support tool that automatically selects the optimal cloud provider and node configuration, considering both cost-effectiveness and runtime efficiency in a multi-cloud, containerized environment.
S16 Unified Template for Microservices Addressing the lack of a unified template and specification for automating microservices deployment across multiple cloud platforms Tool Service Model Tool service model, invocation relationships, service function chains, objective function optimization, accelerated Particle Swarm Optimization, configurable objective parameters in multi-cloud environment
S64 Cloud Interoperability Addressing the interoperability issues between different clouds through automation multi-cloud Harmony Develop an AI-driven multi-cloud management platform to automate cross-cloud interoperability, ensuring seamless data transfer and communication between disparate cloud environments
S63 Cloud Applications Standardization Promoting the adoption of worldwide standards or interfaces for cloud applications and their management through automation Cloud Service Interoperability Cloud service providers and middleware for interoperability, standardized interfaces, and hybrid solutions in multi-cloud environment
S18 Package Creation and Scaling Automating the creation and scaling of packages using VM and Docker based on machine-readable definition files Package Automation with VM Docker Implement a scriptable automation framework that facilitates creation, scaling, and management of packages using VM and Docker based on machine-readable definition files
Category 7: Application and Service Management
S55 5G Ecosystem Component Selection Simplifying the complexity of 5G ecosystem component selection through automation Open-source Software Components Open-source software components for customizable container and virtual machine orchestration in contemporary deployments
S62 End-to-End Tail Latency Management Automating end-to-end tail latency management in a microservice architecture Performance Modeling Approach Modeling approach for performance evaluation of container-level and VM-level data in multi-cloud environment
S79 Distributed Application Complexity Addressing the increasing complexity of managing distributed applications in containerized multi-cloud environment Complexity of Distributed Applications Increasing complexity of distributed applications and operating multi-cloud management topologies involving CMPs
Category 8: Runtime and Service Discovery
S72 Network Complexity Managing the complexity of network control and administration within containerized applications across multiple cloud environments. Network Management Complexity Complexity of network management and control in multi-cloud environment
S83 Social Network Service Discovery Improving service discovery algorithms for tagging in online social networks within containerized multi-cloud environment Automated Discovery and Selection Automating discovery and selection of relevant and appropriate services for composition and QoS management in multi-cloud environment
S08 Complex Application Administration Handling growing application complexity due to new technologies and needs through automation Autonomic Computing Systems Autonomic Computing Systems with knowledge transfer for containerization in multi-cloud environment
[…]
3.7. Deployment Challenges and Solution Framework (RQ4 and RQ5)
Table 13 presents a comprehensive overview of the challenges and corresponding solutions faced during the deployment of containerized applications in multi-cloud environment. As organizations increasingly adopt containerization and multi-cloud strategies to enhance scalability, flexibility, and resource utilization, they encounter various obstacles that require effective solutions. The figure consists of nine categories, each representing a specific deployment challenge, along with the corresponding challenge description and its solution. These categories encompass a wide range of issues, including deployment complexity and Orchestration, access and communication control, security and compliance management, multi-cloud deployment coordination and Integration, monitoring and scalability challenges, microservices architecture and containerization, and network connectivity and hybrid cloud integration challenges.
Deployment Complexity and Orchestration: This category focuses on addressing challenges and providing solutions related to the complexity and orchestration of container deployment in multi-cloud environment. One of the significant challenges faced in this context is Performance Testing Deployment. It involves evaluating testing procedures and assessing performance metrics during container deployment across multiple cloud environments. To overcome this challenge, the corresponding solution is Standardized Benchmarking and Automation. This solution entails adopting Standardized Benchmarking and automatic systems to ensure consistent and reliable performance evaluation. Another challenge is Deployment Validation, which requires conducting thorough validation and artifact checks during container deployment in multi-cloud environment. To effectively address this challenge, the suggested solution is the implementation of a Testing Process Management System. This system enhances and oversees the validation process for deployment, making sure that all essential criteria are met before deployment proceeds. The challenges outlined in this category can simplify and optimise container deployment in complex multi-cloud environment, facilitating efficient and reliable operations across different cloud service providers.
Access and Communication Control: This category emphasizes maintaining secure multi-cloud connectivity and managing user access control. Challenges in this category include ensuring secure access to resources and secure communication between containers and across clouds. The proposed solutions involve implementing encryption and authentication mechanisms, using secure communication protocols, and executing operations in a provider-agnostic manner to maintain security.
Security and Compliance Management in Multi-cloud Deployment: This category focuses on security and compliance challenges faced during multi-cloud container deployment. One challenge, Vendor Lock-in Prevention Deployment, aims to prevent vendor lock-in and ensure modular and loosely-coupled deployment in multi-cloud environment. The solution proposed is the Cloud Modelling Framework, which introduces a framework for specifying provisioning and deployment to enhance compatibility with existing ACSs and cloud solutions. Another challenge in this category is Security Deployment in Multi-container Environment, which addresses security and privacy concerns during multi-container deployment on the same OS kernel in multi-cloud environment. The solution is Cross-Container Isolation, which enforces cross-container isolation to enhance security.
Multi-cloud Deployment Coordination and Integration: This category deals with challenges related to coordination and integration during multi-cloud deployment. One challenge, Multi-cloud PaaS Deployment, aims to overcome high entry barriers for deploying a multi-cloud PaaS infrastructure. The proposed solution is Lightweight Proxies for PaaS Integration, which utilizes lightweight proxies to facilitate individual PaaS services integration. Another challenge is IoT Application Deployment, which addresses challenges related to deploying IoT applications in multi-cloud environment. The solution is based on the DRA Framework and CI Broker for Multi-Cloud Deployment, which leverages the DRA framework and CI broker component for efficient multi-cloud deployment.
Monitoring and Scalability Challenges: This category focuses on challenges related to monitoring, performance, and scalability during multi-cloud container deployment. For instance, Legacy Code Migration Deployment deals with handling legacy code migration, tenant engine separation, and composition modeling during container deployment in multi-cloud environment. The proposed solution is the Reusable Architectural Pattern with Docker and Cloudant, which advocates for adopting a reusable architectural pattern with Docker containers and Cloudant persistence layer.
Microservices Architecture and Containerization: This category addresses challenges associated with microservices architecture and containerization in multi-cloud environment. For example, Autonomic System Deployment aims to implement an autonomic management system for deploying multi-layer, multi-tier web architectures in multi-cloud environment. The solution proposed is the Self-Tuning Performance Model and Autonomic Management, which introduces a self-tuning performance model and autonomic management system.
Network Connectivity and Hybrid Cloud Integration: This category focuses on network connectivity and hybrid cloud integration challenges during multi-cloud deployment. One challenge, Inter-edge Bandwidth Deployment, aims to address challenges related to interconnecting distributed localities and provisioning inter-edge bandwidth.
Cloud Deployment Constraints and Challenges: This category deals with various constraints and challenges encountered during cloud deployment. For instance, ACS Compatibility Deployment aims to ensure compatibility with existing ACSs and cloud solutions during container deployment in multi-cloud environment. The solution proposed is the Scalable Multi-Cloud Deployment Framework, which proposes a scalable deployment framework for multi-cloud applications.
Infrastructure Provisioning and Container Deployment Challenges: This category addresses challenges related to infrastructure provisioning and container deployment in multi-cloud environment. For example, Fog Computing Deployment deals with fog computing hardware and container deployment time challenges. The proposed solution is Reorganized Container Images and Docker Deployment, which involves reorganizing container images and modifying the Docker deployment process.
Application Deployment Workflow and Orchestration Challenges: This category focuses on challenges related to application deployment workflow and orchestration in multi-cloud environment. One challenge is API Management Deployment, which deals with effectively managing different cloud providers’ APIs/interfaces for multi-cloud container deployment. The solution proposed is DevOps Approach for Multi-Cloud Applications, which advocates for adopting a DevOps approach for managing the lifecycle of multi-cloud applications.
Table 13. Deployment challenge and solution framework for containerized applications in multi-cloud environment
ID Challenge Challenge Description Solution Solution Description
Category 1: Deployment Complexity and Orchestration Challenges
S01 Performance Testing Deployment Evaluating testing and performance during container deployment in multi-cloud environment Standardized Benchmarking and Automation Adoption of standardized benchmarking applications and automatic systems
S09 Deployment Validation Conducting rigorous validation and artifact checks during container deployment in multi-cloud environment Testing Process Management System Implementation of a testing process management system
S02 Cost-effective Deployment Ensuring cost-effective and flexible container deployment options in multi-cloud environment Containerization and Microservice Architecture Implementation of containerization and microservice architecture
S38 Resource Management Deployment Managing diverse resources across multiple cloud service providers during container deployment User Preference-Based Multi-Cloud Resource Brokering Adoption of multi-cloud resource brokering dependent on user preferences and application requirements
S50 NFV Site Deployment Managing site selection, dynamic VF allocation, and service chaining during NFV deployment in multi-cloud environment Site Selection and VF Allocation Algorithms Development of algorithms for site selection and dynamic VF allocation
S79 Cloud Management Platform Evaluation Deployment Addressing diverging functionality, lack of standard benchmark, and evaluating CMPs during multi-cloud container deployment Standardized Output Formats and Evaluation Criteria Establishment of standardized output formats and evaluation criteria
S37 Standardized Requirement Deployment Standardizing the expression of application requirements during multi-cloud container deployments Application-Level Resource Managers for Multi-Cloud Emergence of application-level resource managers supporting multi-cloud deployment
S42 Heterogeneous Node Deployment Managing the deployment and control of heterogeneous nodes in multi-cloud environment Technology-Independent Multi-Level Adaptation Utilization of MicroCloud and TOSCA library for technology independence and fine-grained multi-level adaptation
S52 Deployment Plan Generation Identifying suitable cloud services and generating deployment plans for multi-cloud environment Cloud Services Selection and Deployment Planning Utilization of the MUSA Decision Support Tool and Deployer tool for cloud services selection and deployment planning
S57 TOSCA-based Deployment Implementing TOSCA-based container deployment across multiple cloud providers Brokered Multi-Cloud Deployment with TOSCA Adoption of a brokered approach for multi-cloud deployment with TOSCA
S58 Standardization Deployment Addressing compatibility issues and lack of standardization during multi-cloud container deployment Proxy Cloud Virtualization for OpenStack Framework mapping other clouds as sub-clouds to OpenStack with Proxy Cloud Virtualization layer
Category 2: Security and Compliance Management in Multi-Cloud Deployment
S20 Vendor Lock-in Prevention Deployment Developing strategies to prevent vendor lock-in and ensure modular and loosely-coupled deployment in multi-cloud environment Cloud Modelling Framework Introduction of a Cloud Modelling Framework for specifying provisioning and deployment
S25 Security Deployment in Multi-container Environment Addressing security and privacy concerns during multi-container deployment on the same OS kernel in multi-cloud environment Cross-Container Isolation Enforcing cross-container isolation
S34 Security Deployment Ensuring integrated security during the deployment of multi-cloud applications Security Control Specification and Deployment Framework Utilization of the MUSA framework for specifying, deploying, and assessing security controls
S72 Virtual Network Security Deployment Providing secure virtual networks during multi-cloud container deployment Network Virtualization Platform for Multi-Cloud Adoption of Sirius, a network virtualization platform for multi-cloud environment
S83 Malicious Service Management Deployment Managing malicious services and supporting multi-cloud communities formation during container deployment Collusion-Resilient Trust Aggregation Technique Implementation of a trust aggregation technique resilient to collusion attacks
Category 3: Multi-Cloud Deployment Coordination and Integration
S10 Multi-cloud PaaS Deployment Overcoming challenges and high entry barriers for deploying a multi-cloud PaaS infrastructure Lightweight Proxies for PaaS Integration Usage of lightweight proxies for individual PaaS services integration
S85 IoT Application Deployment Addressing challenges related to deploying IoT applications in multi-cloud environment DRA Framework and CI Broker for Multi-Cloud Deployment Utilization of the DRA framework and CI broker component for multi-cloud deployment
Category 4: Monitoring, Performance, and Scalability Challenges
S17 Legacy Code Migration Deployment Handling legacy code migration, tenant engine separation, and composition modeling during container deployment in multi-cloud environment Reusable Architectural Pattern with Docker and Cloudant Adoption of a reusable architectural pattern with Docker containers and Cloudant persistence layer
S23 Tenant Performance Deployment Managing tenant service performance competition during multi-cloud container deployment Multi-Tenant and Multi-Instance Hybrid Deployment Implementation of a multi-tenant and multi-instance hybrid deployment scheme based on container technology
S45 Monitoring Deployment Overcoming the limitations of current monitoring tools during multi-cloud container deployment Platform-Independent Monitoring System Implementation of JCatascopia, a platform-independent and interoperable monitoring system
Category 5: Microservices Architecture and Containerization Challenges
S28 Autonomic System Deployment Implementing an autonomic management system for deploying multi-layer, multi-tier web architectures in multi-cloud environment Self-Tuning Performance Model and Autonomic Management Development of a self-tuning performance model and autonomic management system
S46 Microservice Deployment Managing the complexity of deploying microservice architectures in multi-cloud environment Metrics and Requirements Framework for Microservices Framework for establishing requirements and relevant metrics for the deployment of microservice-based applications
S76 Multi-cloud Application Configuration Deployment Overcoming challenges related to configuring applications during multi-cloud deployment Runtime Environment with Object Store and Artifact Repository Deployment of a runtime environment with an object store service and artifact repository
S80 Elastic Container Platform Dependency Deployment Addressing vendor lock-in and dependency on specific elastic container platforms during multi-cloud deployment Separation of Elastic Platform and Cloud Application Separation of elastic platform definition from cloud application definition
Category 6: Network Connectivity and Hybrid Cloud Integration
S35 Inter-edge Bandwidth Deployment Addressing challenges related to interconnecting distributed localities, provisioning inter-edge bandwidth, and centralized edge nodes during multi-cloud deployment Modular Edge Cloud Computing Architecture Designing a modular edge cloud computing system architecture using containerization techniques
S41 Network Latency Deployment Minimizing the impact of network latency on application performance during multi-cloud deployment VM Selection for Composite Applications Selection of proper VMs from the global cloud marketplace to support multiple composite applications
S43 Latency Reduction Deployment Tackling the issue of high end-to-end latencies during multi-cloud deployment Edge-Cellular Hybrid Infrastructure Provisioning Provisioning infrastructure at the edge of the cellular network combined with traditional cloud infrastructures
S63 Vendor Lock-in Prevention Deployment Preventing negative effects of vendor lock-in during multi-cloud container deployment Interoperability Layer Above Cloud Infrastructure Implementation of a layer above the cloud service provider’s infrastructure for interoperability
Category 7:Cloud Deployment Constraints and Challenges
S08 ACS Compatibility Deployment Ensuring compatibility with existing ACSs and cloud solutions during container deployment in multi-cloud environment Scalable Multi-Cloud Deployment Framework Proposal and development of a scalable deployment framework for multi-cloud applications
S30 Manual Intervention Risk Deployment Managing risks related to manual intervention, vulnerability, and exposed services during multi-cloud deployment Automated Security Measures at Deployment Automated implementation of required security measures at the point of deployment
S44 Multi-provider Deployment Handling the challenges of using multiple providers and technologies during multi-cloud container deployment TOSCA-Based Deployment Modeling Approach Adoption of a TOSCA-based deployment modeling approach
S81 Cloud Provider Transition Deployment Addressing incompatibilities of cloud solution features during multi-cloud deployment Semantic Interoperability in Multi-Clouds Adoption of the FCLOUDS framework for semantic interoperability in multi-clouds
Category 8: Infrastructure Provisioning and Container Deployment Challenges
S32 Fog Computing Deployment Addressing fog computing hardware and container deployment time challenges in multi-cloud environment Reorganized Container Images and Docker Deployment Reorganization of container images and modification of the Docker deployment process
S33 Fog Device Deployment Managing fog device availability and services during multi-cloud container deployment On-Demand Fog and Microservices Deployment Enabling on-demand creation of fogs and deployment of microservices using Docker and Kubeadm
S36 Fog Computing Container Deployment Overcoming low rates of container deployment, especially in fog computing environments FogDocker for Container Deployment Deployment of containers using FogDocker after essential file downloads
S39 Distributed Compute Node Deployment Effectively managing containerized applications across distributed compute nodes in multi-cloud environment Label-Based Scheduling Strategy Development of a label-based scheduling strategy
S40 Kernel Security Deployment Addressing security risks related to kernel sharing, isolation, and OS-level security during multi-cloud container deployment Enhanced Container Security Enhancing container security from the host OS level
Category 9: Application Deployment Workflow and Orchestration Challenges
S27 API Management Deployment Effectively managing different cloud providers’ APIs/interfaces for multi-cloud container deployment DevOps Approach for Multi-Cloud Applications Adoption of a DevOps approach for managing the lifecycle of multi-cloud applications
S29 Web Service Deployment Ensuring correct and efficient deployment of web services across various computing environments Kubernetes-Based Containerized Deployment Deployment of containerized environments using Kubernetes
S51 Complex Application Deployment Overcoming complexities associated with application deployment in multi-cloud environment Aneka Platform for Distributed Applications Adoption of the Aneka platform for programming and deploying distributed applications
S53 High-availability Deployment Ensuring high availability and preventing outages during deployment in multi-cloud environment Middleware Support for High Availability in Multi-Cloud PaaS Introduction of novel Middleware Support for High Availability in multi-cloud PaaS
S69 Complex Workflow Deployment Utilizing multiple cloud providers for deploying complex application workflows CloudSME Simulation Platform Utilization of the CloudSME Simulation Platform for access to various resources and workflow support
S84 Business Process Provisioning Deployment Orchestrating multi-cloud business process provisioning during deployment Multi-Cloud Service Orchestration Frameworks Adoption of multi-cloud service orchestration frameworks
[…]
3.8. Monitoring Challenges and Solution Framework (RQ4 and RQ5)
Table 14 provides an extensive overview of the monitoring challenges and solution framework for containerized applications in multi-cloud environment. It encompasses diverse challenges related to monitoring containerized applications in the complex multi-cloud landscape. Each category in the figure targets specific monitoring aspects and offers solutions to address these challenges.
Performance, Consistency, and Variability Monitoring: This category emphasizes sustaining performance monitoring in dynamically varying cloud environments for containerized applications. It tackles the challenge of ensuring consistent performance monitoring and proposes solutions like performance variation monitoring and benchmarking for optimized performance.
Resource and Infrastructure Monitoring: In this category, the focus is on monitoring resource and infrastructure management for containerized applications in multi-cloud environment. Challenges include network communication, system image and configuration monitoring, and dynamic resource management. Solutions involve service mesh implementation, Docker server deployment, and using Platform-as-a-Service (e.g., Aneka PaaS) for Cloud Computing.
Security, Compliance, and Trust Monitoring: This category addresses the vital aspects of monitoring security, compliance, and trust management for containerized applications in multi-cloud settings. Challenges include accessing provider host information, continuous security control, and multi-cloud security compliance. Solutions include ML-based monitoring and security assurance platforms in the form of a SaaS (e.g.,MUSA), which support multi-cloud application runtime security control and transparency to enhance user trust.
Application Optimization and Adaptability Monitoring: The challenges of this category revolve around monitoring application optimization and adaptability in multi-cloud environment. Challenges include real-time application monitoring, microservices optimization, application behavior, download execution, and design and state monitoring. To address these challenges, solutions encompass the Dynamic Resource Allocation (DRA) framework, Microservices-based Cloud Application-level Dynamic Orchestrator (MiCADO) optimization, Topology and Orchestration Specification for Cloud Applications (TOSCA) event modeling, FogDocker implementation, and adopting a cloud modeling framework.
System Complexity and Standardization Monitoring: This category focuses on addressing system complexity and promoting standardization in monitoring practices across multi-cloud environment. Challenges include proprietary monitoring solutions and resource isolation mechanisms in the Linux kernel. Solutions encompass pervasive monitoring and identifying in-container leakage channels.
Multi-cloud Coordination and Integration Monitoring: In this category, the challenges revolve around coordinating and integrating monitoring data across multiple cloud environments. Challenges include multi-cloud monitoring challenges, cross-level monitoring improvement, and integrated multi-cloud management monitoring. Solutions encompass self-healing techniques, business process monitoring, and integration with tools like Gnocchi.
[…]
Table 14. Monitoring challenge and solution framework for containerized applications in multi-cloud environment
ID Challenge Challenge Description Solution Solution Description
Category 1: Performance Monitoring Consistency and Variability
S01 Performance Monitoring Ensuring sustained performance monitoring in dynamically varying cloud environments Performance Variation Monitoring Monitoring performance fluctuations in containerized applications through benchmarking periods
S02 Benchmarking Performance Variation and Cost Efficiency Monitoring Efficiently tracking performance disparities and cost-effectiveness for refined monitoring in benchmarking SDBO Optimization Maximizing user budget utility with the optimization feature in SDBO
S29 Website Performance and Availability Monitoring Implementing continuous oversight over website availability and performance Kubernetes Monitoring Utilizing Kubernetes for monitoring performance of containerized services in multi-cloud environment
S38 User Experience Variability Monitoring Addressing monitoring fluctuations in user experience based on diverse factors CSP QoS Metrics Collection Collecting Cloud Service Provider (CSP) Quality of Service (QoS) metrics for benchmark data in multi-cloud environment
S43 Independent Performance Metrics Monitoring Establishing application-independent performance metrics for Round Trip Time (RTT) and payload throughput Performance Evaluations Conducting performance evaluations and autonomic resource provisioning for multi-cloud and MEC platforms using containerized applications
S50 Comprehensive Network Performance Measurement Monitoring Ensuring comprehensive performance measurements monitoring and accounting for transmission delays in the network Server Selection Scheme Applying a server selection scheme to minimize latency in multi-cloud environment based on end-to-end performance measurements
S62 Microservice Performance Degradation Monitoring Addressing unpredictable performance degradation in containerized microservices Tail Latency Prediction Implementing a modeling approach to predict end-to-end tail latency in multi-layer cloud data of containerized applications
Category 2: Resource and Infrastructure Management Monitoring
S05 Network Communication Monitoring Managing service-to-service communications over a network Service Mesh Implementation Deploying a network service mesh for service-discovery, secure connectivity, and visibility in multi-cloud platforms
S09 System Image and Configuration Monitoring Monitoring of system images, remote software packages, usage history, and container execution, configuration, and monitoring. Docker Server Deployment Deploying containerized services using Docker Server API version 1.23 in multi-cloud platforms
S21 Vcluster and Framework Management Monitoring Monitoring and managing the performance and resource usage of each Vcluster and application framework can be complex when dealing with multiple tenants and different types of applications. LXC Implementation Using Lightweight Containers (LXC) for efficient resource allocation and scalability in multi-cloud environment
S24 Multi-Cloud Resource Utilization Monitoring Monitoring container performance and resource usage across multiple clouds Prometheus and Grafana Monitoring Implementing Prometheus and Grafana for monitoring performance and availability of containerized applications in multi-cloud environment
S51 Dynamic Resource Management Monitoring Applications should add and release resources dynamically by measuring and monitoring usage Aneka Platform Utilizing the Aneka platform for monitoring resources and managing dynamic resource acquisition and release in multi-cloud environment
Category 3: Security, Compliance, and Trust Management Monitoring
S04 Provider Host Information Access Monitoring Overcoming the restriction in accessing provider host information ML Monitoring Using machine learning to supervise performance of containerized applications and OS usage
S34 Continuous Security Control Monitoring Strengthening continuous monitoring and control over security MUSA Platform Applying the MUSA Security Assurance Platform for continuous monitoring and dynamic security enforcement of containerized applications
S52 Multi-Cloud Security Compliance Monitoring Monitoring security compliance consistently across multiple cloud providers MUSA Security Compliance Using the MUSA Security Assurance Platform to monitor multi-cloud application compliance and ensure security
S53 Service Continuity and Data Privacy Monitoring Ensuring data privacy, uninterrupted service, fault tolerance, and availability NoMISHAP Implementation Implementing the NoMISHAP solution to address monitoring challenges in multi-cloud environment with containerized applications
S83 Trust Relationship and Network Contact Monitoring Building trust relationships, collecting judgments, and maintaining networks of contacts Trust Bootstrapping Implementing a trust bootstrapping mechanism that combines endorsement and decision tree classification for containerized applications in multi-cloud platforms
Category 4: Application Optimization and Adaptability Monitoring
S20 Adaptive System Design and State Monitoring Implementing continuous design with dynamic adaptability; Enhancing system state evaluation through robust monitoring Cloud Modelling Framework Utilizing a Cloud Modelling Framework for continuous design and dynamic adjustment of multi-cloud containerized applications
S36 Partial Download Execution Monitoring Ensuring accurate execution of applications for partially downloaded files FogDocker Implementation Using FogDocker with low runtime overhead to ensure execution correctness of containerized applications
S44 Event-Driven Application Behavior Monitoring Modeling event-triggered behavior of application components TOSCA Event Modeling Applying TOSCA modeling for event sources, events, and functions in containerized applications
S68 Microservice Optimization Monitoring Enhancing surveillance over microservices across multiple cloud environments for optimization MiCADO Optimization Utilizing MiCADO for runtime monitoring and optimization of cloud applications in multi-cloud environment
S85 Real-Time Application Monitoring Real-time monitoring of applications for immediate corrective action DRA Framework Implementing the DRA framework with the Monitoring and Tracking Platform (M4) for containerized applications in multi-cloud platforms
Category 5: System Complexity and Standardization Monitoring
S10 Proprietary Monitoring Solution Standardization Overcoming the constraints of complex, proprietary monitoring solutions and fostering standardization Pervasive Monitoring Implementing pervasive monitoring solutions for containerized applications within multi-cloud PaaS
S25 Kernel Resource Isolation Mechanism Monitoring Addressing the limitations of resource isolation mechanisms in the Linux kernel Leakage Defense Identifying in-container leakage channels and implementing a two-stage defensive approach in multi-cloud environment
S27 Multi-Cloud Application Monitoring Complexity Simplifying monitoring processes in multi-cloud applications DECIDE DevOps Expansion Expanding the DECIDE DevOps framework for native multi-cloud containerized applications
S28 Monitoring Agent Coordination Enhancing the coordination of monitoring agents for improved system behavior Autonomic Management Implementing an autonomic management system with monitoring agents, a performance model, and adaptive actions for containerized applications
S33 Heterogeneous Fog Resource Management Monitoring Efficient management of diverse fog resources Kubernetes Utility Architecture Implementing a master-worker nodes architecture with Kubernetes Utility (Kubeadm) in multi-cloud environment
S42 Application Topology and Dependency Monitoring Enhancing coordinated adaptation through thorough understanding of application topologies and dependencies MicroCloud TOSCA Library Utilizing the MicroCloud TOSCA library for standardized description of topologies in multi-cloud containerized applications
S45 Elastic Cloud Service Monitoring Research Investigating elastic cloud service monitoring as an open research problem JCatascopia Monitoring Utilizing JCatascopia for multi-layer, platform-independent monitoring of cloud containerized applications
S46 Microservice Monitoring Diversity Addressing diverse monitoring requirements and varying metrics quality in microservices QoM Definition Defining Quality of Monitoring (QoM) and implementing an API for cloud provider capabilities in multi-cloud environment
S49 Compute and Network Resource Utilization Monitoring Strengthening surveillance and prediction capabilities of compute and network resource utilization Time Series Resource Utilization Developing algorithms for monitoring and predicting resource utilization of containerized applications using time series data
Category 6: Multi-cloud Coordination and Integration Monitoring
S58 Integrated Multi-Cloud Management Monitoring Implementing integrated monitoring and management across multi-cloud environment Gnocchi Integration Integrating with federated gnocchi for consolidated monitoring and management of multi-cloud environment
S74 Multi-Cloud Monitoring Challenges Addressing monitoring challenges in multi-cloud environment Self-Healing Techniques Implementing a rule-based algorithm for monitoring and self-healing techniques in multi-cloud environment
S84 Cross-Level Monitoring Improvement Ensuring precise and adaptable metrics computation, overcoming inadequate cross-level monitoring approaches Business Process Monitoring Conducting cross-cloud and cross-level business process monitoring and coordinated adaptation of containerized applications
3.9. Tools and Frameworks (RQ6)
In response to our RQ6 concerning the tools and frameworks utilized for developing containerized multi-cloud applications, an in-depth exploration of the selected studies was conducted. This comprehensive analysis led to the identification of a total of 160 distinct tools and frameworks (see Table 15), all of which are systematically classified into 45 overarching categories, further delineated into 6 subcategories. Notably, the three most prominent categories emerged from this investigation are cloud services and management platform, infrastructure management technologies, and networking tools and framework. Each of these categories encapsulates a range of tools aimed at facilitating the seamless development, deployment, and management of multi-cloud applications. In the following we briefly report the each of the category, shedding light on the tools and frameworks that collectively form the foundation for containerized multi-cloud application development.
Cloud Services and Management Platforms: Cloud Services and Management Platforms: This category of tools facilitates the management and orchestration of cloud services across multi-cloud environments. It includes a broad range of platforms covering Infrastructure as a Service (IaaS), Platform as a Service (PaaS), and Software as a Service (SaaS). These tools are essential for provisioning virtual resources, deploying applications, and delivering software across different cloud platforms. Examples include Amazon EC2, Google Cloud App Engine, and Microsoft Office 365.
Infrastructure Management Technologies: This category of tools facilitates the management of cloud infrastructure. It encompasses container platforms, virtualization solutions, Kubernetes services, orchestration platforms, and tools for configuration management and continuous integration. Key technologies like Docker, Kubernetes, Puppet, Chef, and Terraform play a vital role in automating and optimizing the infrastructure underpinning containerized applications.
Monitoring Tools: This category of tools facilitates the monitoring and analysis of cloud and container operations. It includes solutions for cloud monitoring, container monitoring, visualization & alerting, resource monitoring, and logging & analysis. Tools such as Amazon CloudWatch, cAdvisor, Prometheus, Grafana, and the ELK stack are instrumental in ensuring the performance and reliability of cloud services.
Networking Tools: This category of tools facilitates the establishment and management of network functionalities in a multi-cloud environment. It covers load balancers, reverse proxies, network services, controllers, paradigms, types, and identity protocols, along with storage networks and event buses. Technologies like Software-Defined Networking (SDN), Neutron, and Kafka event bus are crucial for maintaining robust and efficient network operations.
Development and Programming: This category of tools facilitates the development and programming aspects in cloud environments. It includes orchestration tools, various programming languages, integrated development environments (IDEs), service meshes, APIs, and cloud modeling tools. Examples are Java, Python, Eclipse IDE, Istio, REST API, and Cloud Modelling Framework (CLOUDMF), which support the creation and deployment of cloud-native applications.
Security and Assurance: This category of tools facilitates the security and compliance of cloud environments. It encompasses vulnerability management, security platforms & tools, access control & policies, Linux security, and cryptography models & tools. Technologies like OpenVAS Vulnerability Scanner, Kubernetes.eBPF, Seccomp-BPF, and Attribute-Based Encryption and Signature are essential for protecting cloud infrastructure and data against various security threats.
Table 15. Tools and frameworks for containerized applications in multi-cloud environment
Subcategory Tools and Frameworks
Category 1: Cloud Services and Management Platforms
Cloud Management Platforms
Cloudyn (S65), IBM Brokerage Services (S65), CloudBroker Platform (S69), CloudcheckR (S79), MistIO (S79), ManageIQ (S67, S79), CMP2 (S79), AWS SDK (S71), CloudSME Simulation Platform (CSSP) (S69), MiCADO (S69), WS-PGRADE/gUSE (S69), Vagrant (S08), AWS OpsWorks (S38), Universal Compute Xchange (UCX) (S65), Open vSwitch (OvS) (S72), Liferay portlets (S69), Cloud Service Providers (CSPs) (S71), AWS Data Pipeline (S78), Cloud computing (S25), Amazon AWS (S42, S43), GENI (S38), Google Cloud Platform (S38, S53), Microsoft Azure (S38, S48, S51, S54, S60), AWS EC2 (S51), IBM (S49), Amazon Web Services (S54), Google Cloud (S54), SalesForce (S53), Multi-Cloud (S60), Cloud computing platforms (S74), Multiple independent clouds (S74), Cloud computing (S84)
Cloud Deployment Model
Public clouds: (S37, S53, S59, S61), Google cloud (S07), Microsoft Azure (S08, S16, S19), AWS (S65, S67, S80, S81), Google’s cloud (S65), Google Compute Engine (S67, S80), Alibaba (S65), Tencent (S65), IBM Watson Health Cloud (S78). Private cloud: OpenStack (S20, S21, S43), VMware vSphere Specialized Cloud Platforms: European Grid Infrastructure Federated Cloud (EGI FC) (S81)
Infrastructure as a Service (IaaS)
Amazon EC2 (S01, S51), AWS (S02, S65, S67, S80, S81), Azure (S02, S08, S16, S19, S67, S80), Google GCE (S03, S67, S80), Aliyun ECS (S03), Baidu BCE (S03), Tencent CVM (S03), JD VM (S03), VMware vSphere (S08), GoGrid (S51), Windows Azure (S51)
Platform as a Service (PaaS)
Google Cloud App Engine (S02, S10, S15), Heroku (S08, S53), IBM Bluemix Workflow Service (S17, s53), Amazon Elastic Beanstalk (S20), Cloud Foundry (S24), OpenShift (S24, S53, S55)
Software as a Service (SaaS)
SaaS (S15), Microsoft Office 365 (S53)
Distributed Processing Frameworks
Hadoop (S21, S37), Spark (S21, S28), Apache Yarn (S37), Storm (S21), Apache Kafka (S34, S35)
Data Storage
Apache Cassandra (S28), Firebase real-time database (S35), Redis (S36, S42), MySQL (S42), NoSQL (S85), SQL (S85)
Messaging Systems
Apache Kafka (S34, S35), Zookeeper (S35)
Category 2: Infrastructure Management Technologies
Container Platforms
Docker (S08, S13, S16, S17, S18, S25, S27, S28, S29, S30, S31, S32, S35, S55, S59, S62, S74, S77, S79, S80), Smart Docker Benchmarking Orchestrator (SDBO) (S01), Docklet (S21), LXC (S24, S25, S29), Docker container engine (S24), Docker container (S23), Docker CE version 18.01 (S36), Docker Hub (S36)
Virtualization
Virtual Machines (S05, S22, S23, S46, S72), Amazon EC2 VMs (S18), Docker container virtualization technology (S74), Xen (S19), VMware (S19), Hyper-V (S19), KVM (S19), OpenVZ (S19), Libvirt API (S19)
Kubernetes Services
Microsoft AKS (S03), Aliyun Kubernetes Engine (S03), Tencent Kubernetes Engine (S03)
Orchestration Platforms
Kubernetes (S18, S22, S29, S55, S56, S59, S62, S63, S80), Docker Swarm (S18, S22, S39, S59, S80), Mesos (S21, S22, S59, S80), Yarn (S21), Kubeadm (S33), Kubernetes for container orchestration (S62), Apache Mesos (S59, S80), MiCADO (S69)
Configuration Management
Puppet (S08), Chef (S08, S30), Terraform (S27, S29), Shell scripts (S44), CloudBees (S08) Continuous Integration CloudBees (S08), Maven (S08), NPM (S08), Github (S30), Github (S30)
Container Isolation
Sandboxie (S24), Warden/Garden (S24), Rocket (S24)
Container Infrastructure
Linux Container (S16), Docker Containers on Linux Host (S13), Kubernetes cluster with a Docker (S04), Amazon AWS ECS (S18), Container Technology (S23)
Category 3: Infrastructure Management Technologies
Load Balancer
Load balancer (S23)
Reverse Proxy
Reverse proxy technology (S23), httpd (S36), nginx (S36)
Ports and Interfaces
TrunkPort (S55)
Network Services
Neutron (S55)
Gateways and Tunnels
Gateways (S64), GRE Tunnel (S72)
Network Controllers
Software-Defined Networking (SDN) controller (S72)
Identity Protocols
Host Identity Protocol (HIP) (S77)
Storage Networks
MCSS (S75), Multi-Cloud storage overlay network (S75)
Network Tools
OpenADN (S50), OpenDayLight (S50), Kuryr (S55), Flannel (S55)
Event Buses
Kafka event bus (S34)
Category 4: Development and Programming
Orchestration
MUSA Deployer tool (S52), MUSA Modeler tool (S52), OSM (S55), ManageIQ (S55)
Programming Languages
Java (S30, S34, S35, S54, S80, S81), Go programming language (S32, S36), Python (S35, S54, S56, S58, S79), C# (S54), Ruby (S54), PHP (S54)
IDEs and Editors
Eclipse IDE with tight integration with OPTIMUS (S27), MCSLA editor (S27)
Service Mesh
Consul (S05), Istio (S05), Istio service mesh (S59)
APIs
REST API (S56), OpenStack API (S64)
Cloud Modeling
Cloud Modelling Framework (CLOUDMF) (S20), Cloud Modelling Language (CLOUDML) (S20, S81), CAMEL modeling language (S52, S80, S84)
Category 5: Development and Programming
Vulnerability Management
OpenVAS Vulnerability Scanner (S30)
Security Platforms and Tools
MUSA Security Assurance Platform (S34), Firewall Deployment (S30)
Access Control and Policies
Access control policies (S25)
Linux Security
Linux top command (S34), Kubernetes.eBPF (S40), Seccomp-BPF (S40), Landlock LSM (S40), Pledge (S40), Security Namespace (S40), Linux Security Modules (LSM) (S40)
Cryptography Models and Tools
Attribute-Based Encryption and Signature (S75), ABEABS cryptographic model (S75)
Category 6: Monitoring Tools
Cloud Monitoring
Amazon CloudWatch (S19)
Container Monitoring
cAdvisor (S22), Prometheus (S22), AlertManager (S22)
Visualization and Alerting
Grafana (S27), Telegraf (S27)
Resource Monitoring
Resource monitoring tool (S60), Ceilometer (S56, S58), Gnocchi (S58)
Logging and Analysis
ELK stack (S74)
[…]
4. Discussion
4.1. Containers in Multi-Cloud: Roles and Strategies (RQ1)
Containers play a crucial role in the development and operation of modern software systems within multi-cloud environment. Understanding the various roles and implementation strategies of containers provides significant insights for both researchers and practitioners.
Multifaceted Role of Containers in Multi-Cloud Environment: Containers fulfill diverse roles within multi-cloud environment, such as application deployment, service hosting, and resource management. They facilitate deployment of applications across a broad range of systems, from web apps to IoT applications. Furthermore, containers contribute to enhancing security in multi-cloud operations. For researchers, this variety of roles signifies a rich area for inquiry, with opportunities to investigate potential security vulnerabilities, management complexities, and performance overheads. For practitioners, understanding these roles can inform efficient and effective deployment strategies, leading to more robust multi-cloud implementations.
Container Technology Features and their Significance: The unique features of container technology, such as lightweight virtualization, compatibility with orchestration tools, and application in edge computing, offer significant advantages. For researchers, these features present opportunities for further exploration, including how these characteristics (e.g., lightweight virtualization) influence container performance and effectiveness in multi-cloud operations. For practitioners, understanding these features can lead to better utilization of container technology and more efficient multi-cloud deployments.
Implementation Strategies and their Real-World Implications: The strategic implementation of containers significantly affects their functionality in multi-cloud environment. Strategies related to deployment, orchestration, security, performance optimization, cloud service management, and data storage can have significant implications. For researchers, these strategies present an area for exploration, including investigation into their real-world implications, benefits, and limitations. For practitioners, a refined understanding of these strategies can lead to more efficient implementations and operations.
Containerization’s Impact on IoT and Application Development: The incorporation of containers into IoT and app development has transform these fields. Containers offer improved software unit encapsulation, portability, and efficiency, which are particularly important in the distributed and resource-constrained nature of IoT. Researchers can further explore potential challenges in detail such as network complexity, security, and the small footprint requirement of IoT devices. For practitioners, understanding the impact of containerization can lead to the development of more robust, efficient, and secure IoT and app solutions.
4.2. Patterns and Strategies (RQ2)
Containerized applications in multi-cloud environment requires an understanding of prevailing patterns and strategies. In this context, we discuss the key findings related identified patterns and strategies presented in Table 9.
Diverse Architectural Patterns and Communication Strategies: Our study found that Microservice Architecture and Service-Oriented Architecture (SOA) are the most common used architectural patterns for designing container-based applications in multi-cloud environment. These results align with the existing literature that emphasizes the use of MSA and SOA because of their support for modular development, scalability, and efficient resource use (Balalaie et al., 2016; Dragoni et al., 2017). The other patterns frequently reported are Service Mesh and Service Chaining, which are employed as communication and networking patterns, likely due to their ability to manage complex, distributed service-to-service communications in a scalable way (Zhamak, 2018). The frequent use of Blue-Green Deployment and Object Store Service deployment patterns has also been identified in this study, which may stem from their ability to minimize downtime and ensure high availability—essential factors in a cloud-based setting (Fowler, 2010). Implications: Practitioners can enhance the design, scalability, and efficiency of their applications in a multi-cloud environment by understanding and applying these architectural and communication patterns. Researchers might explore how these patterns evolve with the emergence of new technologies and standards.
Importance of Container Management and Multi-Cloud Strategies: Our findings report the significance of robust container management and multi-cloud strategies. Techniques like the Linux Container (LXC) project, Docker Container Images, and Lightweight Virtualization are frequently used, emphasizing the industry trend towards containerization due to its benefits in efficiency, scalability, and platform independence (Vaughn, 2020). Furthermore, the popularity of multi-cloud strategies like Hybrid/Multi-cloud Approach and Multi-cloud Load Balancing confirms the growing industry move towards using multiple cloud providers to enhance redundancy, flexibility, and avoid vendor lock-in (Gillam et al., 2013). Implications: Practitioners should focus on gaining expertise in container technologies and multi-cloud strategies to take advantage of their numerous benefits. Meanwhile, researchers might examine the challenges and best practices associated with the deployment and management of container technologies across multiple cloud platforms.
Security, Resilience, and Migration Patterns and Strategies: Our study also identifies patterns and strategies related to security, resilience, and efficient migration strategies in a multi-cloud context. Strategies such as File Encryption, Attribute-Based Encryption, and Secure Data sharing are often employed to ensure data security, an area of growing concern with the increase in cloud-based applications (Svantesson and Clarke, 2010). Moreover, resilience strategies such as Fault-tolerance with Redundant Engines and RAFT Consensus Algorithm are essential to ensure system reliability in distributed, multi-cloud environment (Ongaro and Ousterhout, 2014).
Migration between different cloud environments, facilitated by strategies like Service-oriented Migration and Migration to a Virtualized Container, is another vital aspect, given the growth in cloud services and the need for interoperability (Jamshidi et al., 2013). Migrating applications and data between cloud environments is an important consideration in a multi-cloud scenario (Jamshidi et al., 2013). Strategies such as image-based migration and migration to virtualized containers, like Docker, are commonly used (Vaughn, 2020). These techniques facilitate portability and provide consistent environments across different cloud platforms. Additionally, they have significant implications for scaling and managing applications across different cloud service providers. Implications: These findings suggest that security, resilience, and efficient migration should be key areas of focus for practitioners operating in multi-cloud environment. Researchers may also look into novel techniques for enhancing cloud security, improving system resilience, and easing the process of cloud migration.
4.3. Quality Attributes and Tactics (RQ3)
The exploration and implementation of quality attributes and associated tactics are critical to the successful deployment of containerized applications in multi-cloud environment, shaping performance, security, and compatibility characteristics. In the following, we will discuss the key findings of our study, providing insights into the QAs and tactics specifically tailored for containerized applications operating within multi-cloud environment.
Performance Optimization in Multi-Cloud Environment: Our study identifies the role of machine learning, container technology, and location-aware service brokering in enhancing the performance and efficiency of containerized applications within multi-cloud environment. Hashem et al. previously suggested similar efficiency gains from machine learning and container technology, validating the interpretive value of our results (Hashem et al., 2016). Specifically, the application of machine learning aligns with a broader trend towards intelligent cloud resource management, a domain that Buyya et al. identified as significant in cloud computing research (Buyya et al., 2020). Meanwhile, location-aware service brokering emerges as a novel and impactful strategy, stressing the role of geographical considerations in further optimizing containerized applications’ performance. Implications: For researchers, the results of our study offer opportunities to explore the integration of AI techniques with container-based cloud architectures further, especially the application of location-aware service brokering in multi-cloud environment. For practitioners, the results indicate that a adaptable approach involving AI, container technology, and geographical considerations can lead to significant performance enhancements in multi-cloud environment.
Security Measures in Multi-Cloud Deployments: The results of our study indicates the multi-faceted approach to ensuring security in multi-cloud deployments. It highlights the combination of encryption, network security protocols, machine learning, and user-specified security measures, while also encouraging deployment across different cloud providers. This finding complements Singh et al.’s argument for a comprehensive approach to cloud security, and aligns with the recent concept of distributed cloud security (Singh et al., 2016). Our results also report the importance of user-specified security measures, which are in agreement with the user-centric security model proposed by Shin and Dong-Hee (Shin, 2013). Implications: From a research perspective, the interact between different security measures and their effect on overall system security in multi-cloud environment warrants further investigation. For practitioners, these results point out the importance of employing a comprehensive security approach in multi-cloud deployments, including encryption, network security, machine learning, and user-specific security measures.
Ensuring Compatibility Across Different Cloud Providers: The results of our study suggests strategies to ensure compatibility in multi-cloud environment through standardization of interfaces, componentization, and interoperability approaches. Emphasizing lightweight communication protocols and interoperability between clouds, the results related to compatibility mirrors Mell and Grance’s principles of service-oriented architecture, reinforcing the need for effective communication across different cloud platforms (Mell et al., 2011). Petcu et al. have also previously stressed the importance of standardization and componentization for ensuring compatibility (Petcu et al., 2018). Implications: From a research standpoint, the dynamic between lightweight communication protocols, interface standardization, and overall system compatibility in multi-cloud deployments offers fertile ground for future work. Practitioners, meanwhile, can adopt these approaches to ensure seamless communication and interoperability across different cloud providers.
These findings not only confirm existing literature results but also offer valuable insights into the practical application of machine learning, container technology, and location-aware services in multi-cloud environment. Future research could further explore the intersection of these quality attributes to develop more robust, efficient, and secure multi-cloud applications by using containerise technologies.
4.4. Automation Challenges and Solution Framework (RQ4 - RQ5)
The Automation Challenge-Solution Framework for Containerized Applications in multi-cloud environment (see Table 12) suggests that multi-cloud management for containerized application is a complex task encompassing several dimensions. These include multi-cloud automation, deployment and scaling, resource management, data and application migration, testing and bench marking, standardization and interoperability, application and service management, and runtime and service discovery. This indicates that addressing one challenge may potentially impact the other areas, necessitating a holistic approach to tackle these issues. In the following we discuss the some of the key observations about this framework.
Multi-dimensional Challenges: The Automation Challenge-Solution Framework for Containerized Applications in multi-cloud environment suggests that multi-cloud management is a complex task encompassing several dimensions. These include multi-cloud automation, deployment and scaling, resource management, data and application migration, testing and bench marking, standardization and interoperability, application and service management, and runtime and service discovery. This indicates that addressing one challenge may potentially impact the other areas, necessitating a holistic approach to tackle these issues. The multi-dimensionality of the challenges implies the need for an integrated approach that considers the interdependencies among these dimensions. It emphasizes the need for solutions that cater to this complexity rather than address individual challenges in isolation. Implications: For researchers, these findings underscore the need to develop multi-faceted, holistic solutions that can address the various interconnected challenges in managing multi-cloud environment. For practitioners, this implies the importance of strategic planning and implementing solutions that address the multiple facets of multi-cloud management while considering their potential impact on one another.
Emergence of AI/ML Solutions: The framework highlights the increasing utilization of AI and ML in providing intelligent solutions to manage containerized applications in multi-cloud environment. This suggests the growing maturity and applicability of these technologies in complex scenarios, a significant step towards achieving efficient multi-cloud management. Previous studies (e.g., (de Laat et al., 2020), (Chen et al., 2021), (Gill et al., 2022)) have recognized the potential of AI and ML in resolving complex cloud management issues, which is in line with the findings from the framework. This further validates the significant role AI and ML have started to play in this domain. The increasing adoption of AI/ML in multi-cloud environment signifies an essential evolution in the field. This stresses the need for further research and development in these technologies to exploit their full potential in resolving multi-cloud management complexities. Implications: The growing use of AI/ML solutions implies that researchers need to focus on improving these technologies and adapting them to specific multi-cloud management challenges. For practitioners, this points towards the increasing necessity to invest in AI/ML capabilities and integrate them into their multi-cloud management strategies.
Importance of Standardization and Interoperability: The framework underscores the crucial role of standardization and interoperability in multi-cloud environment, facilitating seamless integration and management across different cloud platforms. Recent literature validates the importance of standardization and interoperability in multi-cloud environment, emphasizing their role in enhancing the efficiency of cloud services (Alonso et al., 2023a). This agrees with our results, further confirming the critical need for these factors in the domain. The need for standardization and interoperability reiterates the necessity for a unified framework or protocol across different cloud environments. It suggests that attention to these aspects can lead to more seamless, efficient, and secure management of multi-cloud services. Implications: For researchers, the results stress the need to investigate methods to achieve better standardization and interoperability in multi-cloud environment. For practitioners, the findings suggest the need to adopt standards and focus on interoperability while selecting and implementing cloud services, ensuring more efficient management across different platforms.
4.5. Security Challenges and Solutions Framework (RQ4 - RQ5)
The framework proposed in our study provides a comprehensive and systematic approach to addressing the complex security issues associated with containerized applications in multi-cloud environment. In the following, we discuss the key takeaway from security challenges and solution framwrork presented in Section 3.5 and Table 11.
Emphasize on Container-Specific Security Mechanisms: Security solutions specifically designed for containerized applications are critical to managing threats in multi-cloud environment. The selected studies suggest a variety of methods, including secure container orchestration, container isolation, and advanced deployment strategies that make use of deep reinforcement learning. This aligns with the existing literature (e.g., (Zhong et al., 2022), (Li et al., 2023) (Combe et al., 2016)) which highlights the unique security considerations brought forth by containerization. Containers, being lighter than traditional virtual machines and offering process-level isolation, have revolutionized the way applications are packaged and run. However, they also introduce new security challenges that need to be addressed by using container-specific security mechanisms. Containerization, while a powerful tool for application deployment, necessitates its own suite of security strategies. Implementing these will be key to leveraging the benefits of containerization in a secure manner. Implications: For researchers, this highlights an area of potential study: the development of advanced container-specific security strategies, including those that use advanced techniques such as machine learning. For practitioners, this suggests the need to be well-versed in the specifics of container security. Leveraging features of containerization platforms, such as Docker’s isolation features, or incorporating tools specifically designed for container security will be crucial.
Implementing Multi-Layered Security and Compliance Measures: Implementing multi-layered security measures that encompass data protection, access control, and communication security, along with compliance with legal and regulatory requirements, is of utmost importance. Solutions range from data encryption, role-based access control, secure container orchestration, adherence to regulations like GDPR (Regulation, 2016). This is consistent with current cybersecurity frameworks, such as the NIST cybersecurity framework (of Standards and Technology, 2018), which emphasize a multi-layered approach to security. Furthermore, several studies point to the criticality of data security, user access control, and secure communication in cloud environments (Austin, 2018) (Singh et al., 2016). In a multi-cloud, containerized environment, it is not enough to secure data; access and communication channels must also be secured, and all actions must comply with legal and regulatory requirements. Implications: For practitioners is to design and implement a comprehensive security strategy that encompasses multiple layers of protection. Compliance with legal and industry standards should be an integral part of this strategy. For researchers, this underlines the need for studies that address the integration of various layers of security and the development of comprehensive security frameworks.
Ensuring Security through Standardization and Interoperability: Ensuring security through standardization and interoperability is highlighted as a critical step. Secure and standardized interfaces, such as OpenStack APIs, can be used to enhance security and performance across multiple cloud providers. This finding is aligned with existing literature, which emphasizes the role of standardization in achieving security and interoperability in multi-cloud environment (Buyya et al., 2010) (Petcu, 2011). By implementing standardized interfaces and practices, organizations can ensure a level of interoperability and security across cloud environments. Implications: For practitioners, this means choosing platforms and tools that support standardized interfaces and practices, or working to implement those standards within their own environments. Researchers, on the other hand, could focus on the development of new standards or the improvement of existing ones to better address security challenges in multi-cloud, containerized environments.
4.6. Deployment Challenges and Solutions Framework (RQ4 - RQ5)
Table 14 offers a comprehensive overview of challenges and their corresponding solutions encountered during the deployment of containerized applications in multi-cloud environment. In the following, we provide a discussion on key findings pertaining to the challenges and solutions related to the deployment of containerized applications in multi-cloud environment.
Standardization and Automation in Deployment: Standardization and automation stand out as foundational pillars for achieving streamlined and uniform deployments in multi-cloud landscapes. As Raj et al. (Raj et al., 2018) highlighted, embracing automated deployment pipelines paired with standardized toolsets can significantly diminish complexity and mitigate human-induced errors within multi-cloud frameworks. The challenges, captured under Performance Testing Deployment, Deployment Validation, and Cloud Management Platform Evaluation Deployment in the table, elucidate the nuanced complexities associated with deployments spanning multiple cloud platforms. Solutions such as Standardized Benchmarking and Automation, Testing Process Management System, and Standardized Output Formats and Evaluation Criteria emphasize the imperative nature of adopting uniform procedures and harnessing automated solutions. Multi-cloud deployments, inherently intricate, benefit profoundly from standardization, bringing forth predictability and coherence, while automation paves the way for diminishing human-driven inconsistencies and bolstering deployment speeds. Implications: For researchers, the quest for optimized frameworks championing standardization and automation within multi-cloud environment presents a promising research trajectory. For practitioners, adeptness with standardized methodologies and tools, combined with proficiency in automation implementation, becomes crucial in navigating the multi-cloud management maze efficiently
Security and Compliance in multi-cloud environment: Security remains paramount in cloud deployments. Fernandes et al. underscore that the labyrinth of diverse platforms and differing standards in multi-cloud configurations amplifies security complexities (Fernandes et al., 2014). Challenges articulated as Security Deployment in Multi-container Environment and Malicious Service Management Deployment in Table 13 accentuate the primacy of security considerations. Propounded solutions like Cross-Container Isolation and Collusion-Resilient Trust Aggregation Technique spotlight the industry’s shift towards advanced and meticulous security protocols within multi-cloud paradigms. The ever-evolving landscape of multi-cloud deployments demands security solutions that are not only rigorous but also malleable and formidable. To attain comprehensive security across diverse platforms, strategies must encapsulate both depth of understanding and breadth of application. Implications: For the academic community, pioneering avant-garde security solutions apt for multi-cloud environment offers a fertile domain of inquiry. Meanwhile, practitioners should enshrine principles of data integrity, trust assurance, and vulnerability mitigation at the heart of their multi-cloud deployment blueprints.
Integration, Monitoring, and Infrastructure Management: As highlighted by Ferrer et al. (Ferrer et al., 2016) the ability to operate seamlessly across disparate cloud providers and maintain flexible infrastructures is paramount in harnessing the full potential of multi-cloud deployments. The presented challenges, including Network Latency Deployment and Vendor Lock-in Prevention Deployment, underscore the indispensability of agility and cross-platform compatibility. Proposed solutions such as VM Selection for Composite Applications and Interoperability Layer Above Cloud Infrastructure accentuate the importance of crafting adaptable infrastructure and formulating strategies to mitigate vendor entrenchment. In the context of multi-cloud environment, flexibility and interoperability transcend mere advantageous attributes; they become pivotal determinants shaping adaptability, scalability, and the sustained success of deployment blueprints. Implications: For the academic community, delving into novel approaches that champion infrastructure pliability and seamless operation across diverse cloud platforms emerges as a prospective research avenue. For industry professionals, anchoring strategies around flexibility and interoperability not only optimizes current deployments but also fortifies them against future technological evolutions and shifting business dynamics.
4.7. Monitoring Challenges and Solutions (RQ4 - RQ5)
Table 14 presents a comprehensive view of the Monitoring Challenge-Solution Framework for Containerized Applications in multi-cloud environment. It covers a wide range of monitoring challenges within the intricate multi-cloud setting. In the following, we are providing discussion on the key findings.
Enhancing Performance Stability through Multi-Cloud Monitoring: In the context of multi-cloud environment, ensuring consistent and reliable performance monitoring for containerized applications can be challenging due to the dynamic nature of cloud resources. Existing studies (e.g., (Zhong et al., 2022) (Mungoli, 2023)) highlighted the impact of varying resource availability on the performance of containerized applications across multiple clouds. They emphasized the need for performance benchmarking to track fluctuations and maintain a consistent user experience. The solutions presented in our framework, such as benchmarking performance variation and utilizing Kubernetes for monitoring, align with existing studies suggestion for benchmarking. Our proposed framework expands on this by incorporating Kubernetes for improved multi-cloud performance monitoring. Our framework also suggest reliable performance monitoring requires a combination of standardized benchmarking techniques and cloud-specific tools like Kubernetes. This approach can effectively address performance variations and ensure a consistent user experience across diverse cloud environments. Implications: For practitioners, adopt benchmarking practices to track performance fluctuations, and leverage Kubernetes for unified multi-cloud performance monitoring. For researchers, explore further refinements in benchmarking methodologies and investigate the impact of dynamic multi-cloud environment on performance consistency.
Streamlining Resource and Network Management across Multi-Clouds: Managing network communications, system images, and resource utilization in a multi-cloud setting poses challenges due to the diversity of cloud platforms. Sedghpour et al. discussed the complexities of managing network communications in multi-cloud environment. They emphasized the need for standardized approaches like service mesh to enhance network visibility and control (Sedghpour and Townend, 2022). Solutions in our proposed framework, such as using Docker Server for containerized services and Prometheus/Grafana for resource monitoring, align with Sedghpour et al.’s (Sedghpour and Townend, 2022) suggestion for standardized approaches. Additionally, our framework extends these solutions to address broader resource management challenges. Employing standardized tools like Docker Server, along with platforms like Prometheus and Grafana, can help streamline resource management and network communication monitoring across diverse cloud platforms. Implications: For practitioners, implement Docker Server and utilize platforms like Prometheus and Grafana for efficient resource monitoring and management across multi-cloud environment. For researchers, investigate further integration possibilities for standardized resource management tools and explore the impact of such tools on overall system performance and reliability.
Strengthening Security, Compliance, and Trust: Ensuring data privacy, continuous security control, and compliance in multi-cloud containerized applications is essential but challenging due to the distributed nature of cloud resources. Existing studies (e.g., (Alonso et al., 2023a) (Raj et al., 2018)) discussed the challenges of ensuring security and compliance across multi-cloud environment. They highlighted the importance of continuous monitoring and enforcement mechanisms to address security and compliance gaps. Solutions presented in our proposed framework, such as using machine learning for performance supervision and employing security assurance platforms, are aligned with existing studies that emphasis on continuous monitoring and enforcement for security and compliance. Implementing advanced techniques like machine learning for monitoring and adopting security assurance platforms can significantly enhance security, compliance, and trust management in multi-cloud containerized applications. Implications: For practitioners, integrate machine learning-based monitoring and adopt security assurance platforms to ensure continuous security control and compliance across diverse cloud environments. For researchers, explore the scalability and effectiveness of machine learning-driven monitoring techniques and investigate novel methods for enforcing security and compliance in multi-cloud setups.
4.8. Tools and Frameworks (RQ6)
This study report 160 distinct tools and frameworks for container based applications in multi-cloud environment, all of which are systematically classified into 6 categories, further delineated into 46 subcategories. In the following we briefly discuss some of the key findings about identified tools and frameworks.
Diverse Cloud Services by Leading Providers: Our study reveals a diverse array of cloud services, including IaaS, PaaS, and SaaS offerings from major providers like AWS, Azure, and Google Cloud, emphasizing the extensive options available to developers (see Table 15). This underscores the significant range of choices available to developers. Noteworthy cloud service providers highlighted in our study include Amazon EC2, Google Cloud App Engine, and Azure, further underscoring this finding. A pivotal insight gleaned from our research underscores the paramount importance of harmonizing cloud service models with specific application prerequisites, thereby facilitating efficient allocation of resources and optimization of costs. Implications: For practitioners, our outcomes underscore the critical necessity of meticulously evaluating and selecting the most fitting service model based on the unique demands of a given project. This strategic approach ensures the best alignment between technology and goals. Moreover, the implications for researchers are promising: an avenue emerges for in-depth exploration of trade-offs and the identification of optimal practices while navigating the intricate array of cloud service models. This exploration holds potential for enhancing both the efficiency and effectiveness of cloud solutions.
Docker and Kubernetes: Pioneers of Cross-Cloud Containerization: Our investigation spotlights the profound significance of Docker and Kubernetes in streamlining the containerization and orchestration of applications across a spectrum of cloud platforms. Docker adeptly encapsulates applications and their dependencies within self-contained containers, while Kubernetes takes center stage as the orchestration juggernaut, automating deployment, scaling, and management of these containers. These technologies receive direct validation through table references, serving as prime exemplars. Our research yields a compelling revelation: Docker and Kubernetes hold pivotal roles in ensuring consistent application deployment and seamless scalability, effectively simplifying the complexities of multi-cloud environment. Implications: Practitioners stand to gain significantly by adopting Docker and Kubernetes, as this potent combination bolsters efficiency and cultivates heightened portability. Meanwhile, researchers have a promising avenue to explore deeper facets, including security, performance optimization, and the intricate domains of advanced use cases for Docker and Kubernetes within multi-cloud landscapes. This exploration promises valuable insights that can finely calibrate the potential of these cutting-edge technologies.
The Rise of DevOps: Embracing Automation Tools: Our exploration spotlights the important role of DevOps practices and tools such as Puppet, Chef, and Terraform within contemporary software development landscapes. This emphasis revolves around the power of automation in facilitating streamlined deployment and management processes. These tools, by enabling consistent provisioning and deployment of infrastructure, effectively curtail errors and ensure the replicability of processes. The table substantiates this notion through references to Puppet, Chef, and Terraform. Our study unearths a pivotal insight: DevOps practices wield the ability to augment the software development life cycle by automating tasks, ushering in a realm of efficient and dependable deployment. Implications: Practitioners are advised to seamlessly integrate DevOps principles into their workflows by harnessing the prowess of these tools. Furthermore, a realm of exploration opens for researchers, inviting deeper analyses into the evolutionary trajectory of DevOps practices and their profound impact on elevating software quality and hastening development velocities.
5. Threats to Validity
This systematic mapping study is susceptible to various threats that could influence its outcomes. To address these potential threats, we adhered to the established guidelines for conducting SMSs and SLRs as outlined in (Keele et al., 2007) and (Petersen et al., 2008b). We further categorized and analyzed the validity threats to our study based on the four distinct types of validity threats mentioned in (Zhou et al., 2016) and (Wohlin et al., 2012). In the subsequent sections, we delve into the specific validity threats that pertain to the different phases of this SMS.
5.1. Internal Validity
Internal validity refers to the factors that could affect the analysis of the data extracted from the selected studies. The threats to internal validity could happen in the following steps of this SMS:
  • Study Search: The potential to overlook studies during the search process requires careful measures. To mitigate this risk, we used a combination of primary and snowballing search methodologies, as detailed in Section 2.2.1. To increase the collection of primary studies during the initial search phase, we also took additional measures to address search-related issues. Specifically, we improved our search strings through pilot searches before applying them to the databases. This process helped to create a more effective search strategy.
  • Study Selection: We have outlined the study screening and selection process in Table 3. To ensure objectivity and eliminate personal bias in study selection, we implemented a two-phase approach: (i) initial study screening, and (ii) qualitative evaluation of the shortlisted studies. During this procedure, the lead two authors conducted the study screening based on the criteria explicitly detailed in Section 3. To assess the objectivity of the screening process and the level of agreement between the two authors, we applied Cohen’s Kappa (Cohen, 1968). In cases where the lead authors did not agree, the second and third authors independently reviewed the disputed studies to reach a consensus. All the researchers involved in this study have extensive knowledge and research experience in containerization applications within a multi-cloud environment.
  • Data Extraction: Bias from researchers during data extraction can pose a significant challenge in both SMSs and SLRs. To address this potential issue, we developed a standardized data extraction form (see Table 5) to ensure consistent data retrieval. Initially, the first and fourth authors extracted the data. If any uncertainties arose regarding the extracted data, comprehensive discussions were convened among all authors. Following the recommendations in (Wohlin et al., 2012), a subset of the extracted data was cross-verified by the second and third authors.
  • Bias on Themes Classification: Misclassification of data and primary studies may result in biases stemming from subjective interpretation. To minimize this risk, we adhered to the thematic analysis guidelines established by Braun et al. (Braun and Clarke, 2006), and implemented a six-step process for thematic classification, as detailed in Section 2.3.
  • Data Synthesis: We employed both qualitative and quantitative approaches to assess the gathered data. Potential biases in data synthesis could influence the interpretation of our findings. To address this concern, we used open coding and constant comparison techniques from Grounded Theory (Glaser and Strauss, 1967) to analyze the qualitative data extracted from the selected studies.
5.2. External Validity
Concerns about external validity relate to the applicability and generalizability of study findings. Our research provides a thorough examination of containerization in multi-cloud environments, with findings, analyses, and conclusions specifically tailored to this domain. To ensure robust external validity, we developed a study protocol that outlines our research methodology. The literature review spanned a decade, from January 2013 to March 2023, and included peer-reviewed articles from eight preeminent databases in the fields of software engineering and computer science, which are listed in Table 2. While the review is anchored in academic research, which may not encapsulate unpublished industry practices, the systematic approach and comprehensive timeframe of our analysis make the insights valuable for both academia and practitioners. To complement and enrich our findings, future work could include an industrial study (e.g., blogs, white papers) that would provide a broader view of the application of containerization in practice
5.3. Construct Validity
Construct validity concerns to the accuracy of the operational measures used for data collection in a study. The primary constructs of this study revolve around two concepts: “containerization” and “multi-cloud environment”. Utilizing imprecise or incomplete search terms, or employing unsuitable search strategies, can lead to potential pitfalls such as overlooking pertinent papers or including numerous irrelevant ones during the search phase, and omitting relevant papers during the selection phase. To counter these risks, we implemented the following measures: (i) we initiated a pilot search to verify the relevance and comprehensiveness of our search terms; and (ii) we searched paper on eight databases renowned for computer science and software engineering research. Additionally, we customized search string according to each database syntax.
5.4. Conclusion Validity
Threats to conclusion validity relate to factors, such as data inaccuracies, that can impede drawing accurate conclusions. To mitigate these threats, we adhered to best practices, including the search protocol, pilot search, and pilot data selection, as recommended by Kitchenham et al. (Keele et al., 2007) and Petersen et al. (Petersen et al., 2008b). Moreover, to further ensure the validity of our conclusions, the authors engaged in multiple brainstorming sessions to collaboratively interpret the results and finalize the conclusions.
6. Related Work
This section reviews the most relevant existing research in terms of secondary studies such as literature surveys, state-of-the-art analysis, systematic reviews, and mapping studies that consolidate published literature on containerization in multi-cloud environments. The review focuses on (i) existing challenges, proposed solutions, and emerging trends detailed in Section 6.1 and (ii) container-based deployment of multi-cloud systems in Section 6.2.
6.1. Challenges, Solutions, and Performance of Multi-cloud Containers
In recent years, several literature surveys have been published to analyze state-of-the-art on the applications (Pahl et al., 2017), tools and technologies (Casalicchio and Iannucci, 2020), potential and limitations (Bentaleb et al., 2022), as well as emerging trends (Watada et al., 2019) of container-based solutions for cloud computing systems, discussed below.
6.1.1. Challenges, Applications, Tools, and Emerging Trends
Containers in multi-cloud environments provide a standardized and portable way to package, deploy, and run applications, ensuring seamless deployment and management across diverse cloud platforms (Alonso et al., 2023b) (Saxena et al., 2021). One of the earliest SLR-based studies by Pahl et al. (Pahl et al., 2017) on containerized clouds reviewed a total of 46 studies to identify, taxonomically classify and systematically compare the existing research on containers and their application in the context of cloud-based systems. The SLR classified and compared the selected research studies using a conceptual framework to highlight existing trends and needs for future research. SLR results indicate container orchestration, microservice delivery, continuous development and deployment as emerging trends of research on containerized clouds. The SLR in (Alonso et al., 2023b) reviewed a total of 88 studies (published from 2011 - 2021) on multi-cloud systems. Similar survey studies such as (Saxena et al., 2021) concepts, challenges, requirements and future directions for multi-cloud environments are discussed. A survey of existing approaches and solutions provided by different multi-cloud architectures is entailed along with analysis of the pros and cons of different architectures while comparing the same (Imran et al., 2020).
6.1.2. Performance and Resource Utilization
In a study by Bentaleb et al. (Bentaleb et al., 2022), the authors follow the taxonomical classification from (Pahl et al., 2017) to categorize container-based technologies for cloud systems. The study argues for the need of performance metrics to objectively define and evaluate quality attributes such as resource virtualization, service elasticity, orchestration, and multi-tenancy in containerized cloud systems. The study also highlights the needs for future research in terms of best practices, i.e., patterns and tactics (enabling reuse) and tools (supporting automation) for container-based cloud deployment. In the context of performance, the studies (Casalicchio and Iannucci, 2020) (Watada et al., 2019) report literature review and experimental analysis of factors that influence performance of container-based cloud systems. Specifically, Casalicchio et al. (Casalicchio and Iannucci, 2020) reviewed a total of 97 research studies investigating performance evaluation and run-time adaptation of container-based solution. The study highlighted several unsolved challenges such as I/O throughput optimization, performance prediction, and multi-layer monitoring performance bottlenecks. Moreover, Watada et al. (Watada et al., 2019) conducted an experimental study to compare the performance of VMs, containers and uni kernels in terms of and technological maturity using standard benchmarks and observed containers to optimise performance of container-based. The performance of containerized clouds is also determined by resource utilization, which includes but is not limited to CPU utilization, memory footprints, energy consumption, network bandwidth, and execution time.
Kapil et al. (Netaji and Bhole, 2021) reviewed 64 research papers to gain insight into resource allocation, management, and scheduling. Furthermore, limitations of existing resource allocation algorithms are discussed, indicating the need to investigate algorithms or techniques of performance optimization of containers for cloud systems. The study by Maenhaut et al. (Maenhaut et al., 2020) provided an overview of the current state of the art regarding resource management within the broad sense of cloud computing, complementary to existing surveys in the literature. The study investigated how research is adapting to the recent evolution within cloud solutions, including container technologies.
6.2. Container-based Cloud Orchestration, Deployment, and Security
This section reviews the most relevant research on the container-based deployment with a focus on orchestration and security issues of cloud-based systems, detailed below.
6.2.1. Orchestration and Deployment
Naweiluo et al. (Zhou et al., 2022) explored containerization in High Performance Computing (HPC) environments, contrasting it with cloud computing. The results of this study indicate that containers enhance application deployment efficiency, but face challenges in HPC due to high security levels and the need for extensive libraries and packages that affect cloud portability, often resulting in vendor lock-ins. The study also provides a survey and taxonomy on containerization and orchestration efforts in HPC, pointing out differences with cloud environments and identifying future research and engineering potentials. A systematic mapping study by Naylor et al.(Bachiega et al., 2018) explores virtualization in container-based cloud deployments. The mapping study, based on a comprehensive review of major databases, identifies a significant research gap in the performance evaluation of containers, underscoring the need for further investigation on cloud deployment. Carlos et al. (Awada, 2018) present an optimization strategy for deploying microservices in multi-cloud environments focusing on minimizing cloud service costs, network latency, and startup time for new microservices. Their approach uses a Non-dominated Sorting Genetic Algorithm II (NSGA-II) compared to a Greedy First-Fit algorithm, demonstrating a 300% improvement with NSGA-II. This highlights its effectiveness in container and VM orchestration, offering a significant enhancement in deployment of multi-cloud solution in containers. Emiliano et al. (Paladi et al., 2018) survey container orchestration, proposing a reference architecture for autonomic orchestrators, and identifying research challenges in the field. Their work emphasizes the importance of container technologies in cloud environments and the need for advanced orchestration solutions to manage complex multi-container applications effectively. Uchechukwu Awada (Guerrero et al., 2018a) reviews container orchestration tools and platforms, comparing the architectures, components, and capabilities of several Container Service Platforms (CSPs) and orchestration tools like Amazon ECS, Kubernetes, Docker Swarm, and Mesos. The study offers insights into the current state of container orchestration and suggests future research directions, serving as a guide for developers and organizations.
Koustabh et al. (Dolui and Kiraly, 2018) explore multi-container deployment on IoT gateways to meet the stringent latency requirements of advanced IoT applications. Through their study within the AGILE project, they highlight containerization’s role in overcoming the diversity and resource constraints of IoT gateways, showcasing containerization’s advantages for IoT gateway performance optimization. Their research underscores the potential of containerized environments in enhancing application compatibility, portability, and efficient deployment across diverse hardware architectures. Matteo et al. (Nardelli et al., 2018) introduce the Adaptive Container Deployment (ACD) model to optimize containerized application deployment in geo-distributed environments, focusing on IoT and fog computing resources. ACD, formulated as an Integer Linear Programming problem, aims to improve application performance by leveraging containers’ horizontal and vertical elasticity. The study evaluates ACD’s effectiveness against greedy heuristics, highlighting the need for advanced orchestration solutions to exploit emerging computing environments’ characteristics efficiently.
6.2.2. Security of Containerized Clouds
Nicolae et al. (Casalicchio, 2019) analyze security challenges in cloud orchestration for multi-cloud deployments, proposing a security-enabled orchestration framework. Their research identifies potential attack scenarios and security enforcement mechanisms, aiming to enhance security guarantees for cloud operators in a multi-cloud setting. Mohammad et al. (Bokhari et al., 2018) offer a detailed analysis of cloud computing, defining its essential characteristics, architecture, service models (SaaS, PaaS, IaaS), and deployment models. They discuss the security requirements for public and private clouds across different service models, aiming to provide researchers with a comprehensive understanding of cloud computing’s potential and security challenges. Sari et al. (Sultan et al., 2019) survey the landscape of container security, addressing challenges and solutions across four generalized use cases within the host-container threat landscape. Their analysis covers both software-based solutions utilizing Linux kernel features and hardware-based solutions for enhanced security. The study aims to clarify container security requirements and encourage further research in addressing potential vulnerabilities and attacks. Hendrik et al. (Graupner et al., 2015a) tackle privacy, security, and trust issues in cloud computing through a multi-cloud architecture perspective. They propose a novel technique for enhancing security and unifying access control mechanisms across cloud providers, addressing the challenges of inadequate cross-provider APIs and non-unified access control. Rohan et al. (Gupta et al., 2016) propose a security-enhancing methodology for cloud data storage using container clustering and Docker instances for on-demand encryption. This approach aims to secure data while optimizing resource usage and cost, highlighting the benefits of container technology in improving cloud computing security and efficiency.
6.3. Conclusive Summary
Current survey-based studies and systematic reviews, discussed above (e.g., (Casalicchio and Iannucci, 2020) (Pahl et al., 2017)), aim to consolidate the latest findings in published research concerning container-based solutions for cloud computing systems. These studies investigated various aspects of multi-cloud, including exploring applications, tools, and technologies, assessing potential and limitations, and identifying emerging trends.
Our study distinguishes itself from prior research in several ways. First, our work systematically categorizes container utilization in multi-cloud environments into coherent categories and themes. Some of the key themes are ‘Scalability and High Availability’, ‘Performance and Optimization’, ‘Security and Privacy’, and ‘Multi-Cloud Container Monitoring and Adaptation’. We identified and classified seventy-four patterns and strategies for containerization, which existing studies have not provided. Furthermore, we explored the QAs and associated tactics for containerization in multi-cloud environment. This contribution offers practitioners actionable insights into tailoring their containerization strategies to meet desired quality standards. Lastly, our study introduces four distinct frameworks consisting of challenges and solutions related to security, automation, deployment, and monitoring within the context of containerization in multi-cloud environment.
7. Conclusions
This SMS presents the current state of research regarding containerization in multi-cloud environment, focusing on aims of selected studies, the roles of containers, containers implementation strategies, architectural patterns and strategies, quality attributes and corresponding tactics. Additionally, this SMS also explores deployment, monitoring, security challenges, accompanied by their respective solutions, tools, and frameworks used to implement containerized applications in multi-cloud environment. From an initial pool of extensive search results, 86 relevant studies are were finally chosen to answer the research questions presented in Table 1. The key findings of this SMS are listed below.
  • The findings of this SMS reveal various insights regarding SMS demographics, publishers, publication types, authors’ affiliations, and research themes. The yearly distribution reached its peak in 2016, signifying a shifting level of interest over time. The field of publishing is largely dominated by IEEE (72.09%), with conferences emerging as the primary avenue for publication. Notably, academia (93.02%) stands out as the leading affiliation among authors. The prominent research themes encompass Orchestration and Management, Scalability and High Availability, Performance and Optimization, and Security and Privacy. These themes underscore crucial aspects such as container management, scalability, performance, and security within the context of multi-cloud environment.
  • This SMS identifies and classifies a total of 74 patterns and strategies across 10 subcategories and 4 categories for container-based applications in multi-cloud environment. These patterns and strategies encompass a wide range of aspects such as cloud architecture models, communication and networking, deployment, service models, cloud management, container management, edge computing, IoT strategies, security, resilience, fault-tolerance, and cloud migration. These findings are significant for the application and dvancement of containerization-based applications in multi-cloud environment, providing valuable insights into optimizing architecture, communication, management, security, and migration strategies.
  • This SMS identifies and classifies QAs and tactics from selected studies regarding containerized applications in multi-cloud environment. A total of ten QAs and their related terms are identified. Along with these QAs, a total of 47 tactics are identified to enhance the QAs. Notable takeaways include the utilization of machine learning and location-aware service brokering to enhance the performance and efficiency of container-based applications in multi-cloud environment.
  • This SMS presents a comprehensive security challenge-solution framework for containerized applications in multi-cloud environment. This framework offers valuable insights and actionable strategies for practitioners and researchers to enhance the security of containerized applications, address complex multi-cloud security concerns, and ensure compliance with various regulatory requirements, thereby enabling the development of more robust and secure containerized applications in multi-cloud environment.
  • This SMS presents a framework for the automation challenge-solution framework for containerized applications in multi-cloud environment. The automation challenge-solution framework offers a comprehensive guide for practitioners and researchers to address the complex challenges associated with automating various aspects of deploying, managing, scaling, testing, migrating, and ensuring interoperability of containerized applications in multi-cloud environment. By offering specific solutions for each challenge, this framework aids in the development of more efficient, secure, and adaptable automation processes, empowering practitioners to streamline their operations and enabling researchers to advance the field of multi-cloud containerization.
  • The third framework presented in this SMS is the deployment challenge-solution framework for containerized applications in multi-cloud environment. This framework provides valuable insights and practical guidance to practitioners and researchers alike, aiding practitioners in navigating the complexities of multi-cloud deployment through specific solutions to challenges, while also serving as a comprehensive resource for researchers to delve into various aspects of containerized application deployment, orchestration, security, scalability, and more, thus contributing to the advancement of knowledge and innovation in the field of multi-cloud deployment.
  • Fourth framework presented in this SMS is monitoring challenge-solution framework for containerized applications in multi-cloud environment. This framework serves as a comprehensive guide for practitioners, offering specific solutions to challenges related to monitoring performance, resource management, security, optimization, system complexity, and multi-cloud coordination, facilitating their understanding of complex monitoring scenarios and providing actionable insights. Furthermore, researchers can benefit from this framework by gaining a deeper insight into the nuances of monitoring containerized applications in multi-cloud environment, and it can serve as a foundation for further research and innovation in the field of multi-cloud monitoring.
The findings of this SMS will benefit researchers who are interested in understanding the state of research on containerized applications in multi-cloud environment and conducting further investigations to address the open research issues highlighted in Section 4. Additionally, the insights gained from this SMS will support knowledge transfer to practitioners by providing insights into the challenges, solutions, and methods for monitoring, securing, and optimizing containerized applications in multi-cloud environment. We emphasize the importance for practitioners to develop targeted solutions that effectively tackle monitoring, security, and performance degradation concerns within multi-cloud environment deployment. As a future endeavor, we intend to enhance our SMS by conducting industrial case studies with companies, thereby obtaining practical insights and perspectives from practitioners on the effectiveness and applicability of our proposed frameworks in real-world scenarios. This approach will allow us to bridge the gap between research and practical implementations concerning containerized applications in multi-cloud environment and contribute to the advancement of both academic and industry understanding in this domain.
Acknowledgments
This work is partially sponsored by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311.
8. Appendices
List of selected studies for containerization in multi-cloud environments
Study ID Title Venue Year Reference
[…]
Study ID Title Venue Year Reference
[…]
A component-based adaptation approach for multi-cloud applications
IEEE Conference on Computer Communications (INFOCOM) Workshops
2014 (Almeida et al., 2014)
S02
A Cost-Efficient Multi-cloud Orchestrator for Benchmarking Containerized Web-Applications
International Conference on Web Information Systems Engineering (WISE)
2019 (Jha et al., 2019)
S03
A framework to support multi-cloud collaboration
IEEE World Congress on Services (SERVICES)
2020 (Hua et al., 2020)
S04
A Machine Learning Model for Detection of Docker-based APP Overbooking on Kubernetes
International Conference on Communications (ICC)
2021 (Ramos et al., 2021)
S05
A multi-cloud service mesh approach applied to Internet of Things
Annual Conference of Industrial Electronics Society (IECON)
2022 (Gattobigio et al., 2022)
S06
A Secure Container Placement Strategy Using Deep Reinforcement Learning in Cloud
International Conference on Computer Supported Cooperative Work in Design (CSCWD)
2022 (Deng et al., 2022)
S07
A Service Driver Based Application Execution and Development Method in Multi-cloud Context
International Conference on Data Science and Computer Application (ICDSCA)
2021 (Zhang et al., 2021)
S08
An adaptable framework to deploy complex applications onto multi-cloud platforms
International Conference on Computing Communication Technologies - Research, Innovation, and Vision for Future (RIVF)
2015 (Pham et al., 2015)
S09
An agile container-based approach to TaaS
International Congress Future Telecommunications: Infrastructure and Sustainability (FITCE)
2016 (Verdugo et al., 2017)
S10
An Efficient and Reliable Multi-Cloud Provider Monitoring Solution
Conference on Global Communications (GLOBECOM)
2022 (Sabbioni et al., 2020)
S11
An Efficient Multi-Keyword Search Scheme over Encrypted Data in Multi-Cloud Environment
International Conference on Smart Cloud (SmartCloud)
2015 (He et al., 2022)
S12
An Empirical Investigation of a Fault Tolerant Containerized Application Deployment
International Conference on Informatics (ICI)
2015 (Bisht and Kaur, 2022)
S13
Applying Computational Intelligence for enhancing the dependability of multi-cloud systems using Docker Swarm
Symposium Series on Computational Intelligence (SSCI)
2017 (Naik, 2016a)
S14
Automatically Design Heuristics for Multi-Objective Location-Aware Service Brokering in Multi-Cloud
International Conference on Services Computing (SCC)
2020 (Chen et al., 2022)
S15
Autonomic fine-grained migration and replication of component-based applications across multi-clouds
National Foundation for Science and Technology Development Conference on Information and Computer Science (NICS)
2022 (Pham and Pham, 2015)
S16
Band-Area Resource Management Platform and Accelerated Particle Swarm Optimization Alg. for Container Deployment
IEEE Access
2016 (Ouyang et al., 2022)
S17
Building a Multi-tenant Cloud Service from Legacy Code with Docker Containers
International Conference on Cloud Engineering (IC2E)
2022 (Slominski et al., 2015)
S18
Building cloud native application - Analysis for multi-component application deployment
International Conference on Computer Communication and Informatics (ICCCI)
2021 (Pai and Kumar, 2021)
S19
CLAMS: Cross-layer Multi-cloud Application Monitoring-as-a-Service Framework
International Conference on Services Computing (SCC)
2014 (Alhamazani et al., 2014)
S20
Cloud MF: Applying MDE to Tame the Complexity of Managing Multi-Cloud Applications
International Conference on Utility and Cloud Computing (UCC)
2014 (Ferry et al., 2014)
S21
Cluster as a Service: A Container Based Cluster Sharing Approach with Multi-user Support
IEEE Symposium on Service-Oriented System Engineering (SOSE)
2016 (Cui et al., 2016)
S22
Container Deployment Strategy for Edge Networking
Workshop on Middleware for Edge Clouds and Cloudlets (MECC)
2019 (Wong et al., 2019)
S23
Container-Based Performance Isolation for Multi-Tenant SaaS Applications in Micro-Service Architecture
Journal of Physics: Conference Series
2020 (Wang et al., 2020)
S24
Containerization and the PaaS Cloud
IEEE Cloud Computing
2015 (Pahl, 2015)
S25
ContainerLeaks: Emerging Security Threats of Information Leakages in Container Clouds
International Conference on Dependable Systems and Networks (DSN)
2017 (Gao et al., 2017)
S26
ContMEC: An Architecture of Multi-access Edge Computing for Offloading Container-Based Mobile Applications
IEEE International Conference on Communications (ICC)
2022 (Watanabe et al., 2022)
S27
DECIDE: An Extended DevOps Framework for Multi-Cloud Applications
International Conference on Cloud and Big Data Computing (ICCBDC)
2022 (Alonso et al., 2019)
S28
Delivering Elastic Containerized Cloud Applications to Enable DevOps
International Workshop on Software Engineering for Adaptive and Self-Managing Systems (SEAMS)
2017 (Barna et al., 2017)
S29
Deployment of Multi-Tier Application on Cloud and Continuous Monitoring using Kubernetes
International Conference on Electrical, Electronics, Communication, Computer Technologies and Optimization Techniques (ICEECCOT)
2021 (Gupta et al., 2021)
S30
Deployment-Time Multi-Cloud Application Security
International Conference on Smart Computing (SMARTCOMP)
2017 (Sheridan et al., 2017)
S31
DEVSML 3.0 Stack: Rapid deployment of DEVS Farm in distributed cloud environment using microservices and containers
Symposium on Theory of Modeling and Simulation (TMS/DEVS)
2017 (Mittal and Risco-Martín, 2017)
S32
Disaster Recovery and Business Continuity for Database Services in Multi-Cloud
International Conference on Computer Applications & Information Security (ICCAIS)
2018 (Al-shammari and Alwan, 2018)
S33
Dynamic On-Demand Fog Formation Offering On-the-Fly IoT Service Deployment
IEEE Transactions on Network and Service Management
2020 (Sami and Mourad, 2020)
S34
Dynamic security assurance in multi-cloud DevOps
IEEE Conference on Communications and Network Security (CNS)
2017 (Rios et al., 2017)
S35
Engineering and Experimentally Benchmarking a Container-based Edge Computing System
International Conference on Communications (ICC)
2020 (Carpio et al., 2020)
S36
FogDocker: Start Container Now, Fetch Image Later
International Conference on Utility and Cloud Computing (UCC)
2019 (Lorenzo et al., 2019)
S37
Importance of Application-Level Resource Management in Multi-Cloud Deployments
International Conference on Cloud Engineering (IC2E)
2019 (Dimitrijevic et al., 2019)
S38
Knowledge-Engineered Multi-Cloud Resource Brokering for Application Workflow Optimization
IEEE Transactions on Network and Service Management
2022 (Pandey et al., 2022)
S39
Label-affinity-Scheduler: Considering Business Requirements in Container Scheduling for Multi-Cloud and Multi-Tenant
Brazilian Symposium on Computing System Engineering (SBESC)
2022 (Altran et al., 2022)
S40
Leveraging kernel security mechanisms to improve container security: A survey
International Conference on Availability, Reliability and Security (ARES)
2019 (Bélair et al., 2019)
S41
Location-Aware and Budget-Constrained Service Deployment for Composite Applications in Multi-Cloud Environment
IEEE Transactions on Parallel and Distributed Systems
2020 (Shi et al., 2020)
S42
MicroCloud: A Container-Based Solution for Efficient Resource Management in the Cloud
International Conference on Smart Cloud (SmartCloud)
2016 (Baresi et al., 2016)
S43
Mobile Edge as Part of the Multi-Cloud Ecosystem: A Performance Study
Euromicro International Conference on Parallel, Distributed and Network-Based Processing (PDP)
2019 (Dreibholz et al., 2019)
S44
Modeling and Automated Deployment of Serverless Applications Using TOSCA
Conference on Service-Oriented Computing and Applications (SOCA)
2018 (Wurster et al., 2018)
S45
Monitoring Elastically Adaptive Multi-Cloud Services
IEEE Transactions on Cloud Computing
2018 (Trihinas et al., 2015)
S46
Monitoring-Aware Optimal Deployment for Applications Based on Microservices
IEEE Transactions on Services Computing
2021 (Fadda et al., 2019)
S47
MSSF: A Step towards User-Friendly Multi-cloud Data Dispersal
International Conference on Cloud Computing (CLOUD)
2014 (Libardi et al., 2014)
S48
Multi-agent QoS-aware autonomic resource provisioning framework for elastic BPM in containerized multi-cloud environment
Journal of Ambient Intelligence and Humanized Computing
2022 (Saif et al., 2023)
S49
Multicloud- auto scale with prediction and delta correction algorithm
International Conference on Intelligent Computing, Instrumentation and Control Technologies (ICICICT)
2019 (Harwalkar et al., 2019b)
50
Multi-cloud Distribution of Virtual Functions and Dynamic Service Deployment: Open ADN Perspective
IEEE International Conference on Cloud Engineering (IC2E)
2015 (Bhamare et al., 2015)
S51
Multi-cloud resource provisioning with Aneka: A unified and integrated utilization of Microsoft Azure and Amazon EC2
International Conference on Computing and Network Communications (CoCoNet)
2015 (Buyya and Barreto, 2015)
S52
MUSA Deployer: Deployment of Multi-cloud Applications
IEEE International Workshop on Enabling Technologies: Infrastructure for Collaborative Enterprises (WETICE)
2017 (Casola et al., 2017)
S53
NoMISHAP: A Novel Middleware Support for High Availability in Multicloud PaaS
IEEE Cloud Computing
2017 (Acquaviva et al., 2017)
S54
Object storage in the cloud and multi-cloud: State of the art and the research challenges
IEEE International Conference on Automation, Quality and Testing, Robotics (AQTR)
2018 (Bucur et al., 2018)
S55
On the Deployment of an Open-Source, 5G-Aware Evaluation Testbed
International Conference on Mobile Cloud Computing, Services, and Engineering (MobileCloud)
2018 (Bolivar et al., 2018)
S56
OpenStack Federation in Experimentation Multi-cloud Testbeds
IEEE International Conference on Cloud Computing Technology and Science (CloudCom)
2013 (del Castillo et al., 2013)
S57
Orchestrated Multi-Cloud Application Deployment in OpenStack with TOSCA
International Conference on Smart Computing (SMARTCOMP)
2017 (Tricomi et al., 2017)
S58
Orchestration Based Hybrid or Multi Clouds and Interoperability Standardization
IEEE International Conference on Cloud Computing in Emerging Markets (CCEM)
2018 (Sitaram et al., 2018)
S59
Performance Analysis of Zero-Trust multi-cloud
IEEE International Conference on Cloud Computing (CLOUD)
2021 (Rodigari et al., 2021)
S60
Policies Based Container Migration Using Cross-Cloud Management Platform
International Conference on Information and Automation for Sustainability (ICIAfS)
2018 (Janarthanan et al., 2018)
S61
Policy-based deployment in a hybrid and multi-cloud environment
International Conference on Cloud Computing and Services Science (CLOSER)
2019 (Di Modica et al., 2019)
S62
Predicting the End-to-End Tail Latency of Containerized Microservices in the Cloud
International Conference on Cloud Engineering (IC2E)
2019 (Rahman and Lama, 2019)
S63
Preventing vendor lock-ins via an interoperable multi-cloud deployment approach
International Conference for Internet Technology and Secured Transactions (ICITST)
2017 (Pellegrini et al., 2017)
S64
Private STaaS with OpenStack Cinder Volumes for Hybrid/Multi-cloud
International Conference on Advances in Computing and Communication Engineering (ICACCE)
2019 (Harwalkar et al., 2019a)
S65
Real-World Experience with a Multicloud Exchange
IEEE Cloud Computing
2017 (Zeck and Bouroudjian, 2017)
S66
Reliability Management for Blockchain-Based Decentralized Multi-Cloud
International Symposium on Cluster Computing and the Grid (CCGRID)
2020 (Aral et al., 2020)
S67
Reliability-Aware Cost-Efficient Scientific Workflows Scheduling Strategy on Multi-Cloud Systems
IEEE Transactions on Cloud Computing
2022 (Tang, 2021)
S68
Resource optimization of container orchestration: a case study in multi-cloud microservices-based applications
The Journal of Supercomputing
2018 (Guerrero et al., 2018b)
S69
Scalable multi-cloud platform to support industry and scientific applications
International Convention on Information and Communication Technology, Electronics and Microelectronics (MIPRO)
2018 (Kiss, 2018)
S70
Search-based Methods for Multi-Cloud Configuration
International Conference on Cloud Computing (CLOUD)
2022 (Lazuka et al., 2022)
S71
Secure access control for multi-cloud resources
IEEE Conference on ALocal Computer Networks Workshops (LCN Workshops)
2015 (Graupner et al., 2015b)
S72
Secure and Dependable Multi-Cloud Network Virtualization
International Workshop on Security and Dependability of Multi-Domain Infrastructures
2017 (Alaluna et al., 2017)
S73
Secure and Parallel Expressive Search over Encrypted Data with Access Control in Multi-CloudIoT
Cloudification of the Internet of Things (CIoT)
2018 (Guechi and Maamri, [n. d.])
S74
Self-healing cloud services in private multi-clouds
International Conference on High Performance Computing & Simulation (HPCS)
2018 (Mfula and Nurminen, 2018)
S75
SpyStorage: A Highly Reliable Multi-Cloud Storage with Secure and Anonymous Data Sharing
International Conference on Networking, Architecture, and Storage (NAS)
2017 (Shen et al., 2017)
S76
Support Services for Applications Execution in Multi-clouds Environments
International Conference on Autonomic Computing (ICAC)
2016 (Pop et al., 2016)
S77
SynAPTIC: Secure and persistent connectivity for containers
International Symposium on Cluster, Cloud and Grid Computing (CCGRID)
2017 (Ranjbar et al., 2017)
S78
System Restore in a Multi-cloud Data Pipeline Platform
International Conference on Dependable Systems and Networks – Industry Track
2019 (Wang et al., 2019)
S79
Systematic and Recomputable Comparison of Multi-cloud Management Platforms
International Conference on Cloud Computing Technology and Science (CloudCom)
2018 (Serhiienko and Spillner, 2018)
S80
Towards a lightweight multi-cloud DSL for elastic and transferable cloud-native applications
International Conference on Cloud Computing and Services Science (CLOSER)
2018 (Quint and Kratzke, 2018)
S81
Towards Formal-Based Semantic Interoperability in Multi-Clouds: The FCLOUDS Framework
IEEE International Conference on Cloud Computing (CLOUD)
2017 (Challita et al., 2017)
S82
Towards Model-Driven Provisioning, Deployment, Monitoring, and Adaptation of Multi-cloud Systems
IEEE International Conference on Cloud Computing (CLOUD)
2013 (Ferry et al., 2013)
S83
Towards Trustworthy Multi-Cloud Services Communities: A Trust-Based Hedonic Coalitional Game
IEEE Transactions on Services Computing
2018 (Wahab et al., 2016)
S84
Multi-cloud provisioning of business processes
Journal of Cloud Computing
2019 (Kritikos et al., 2019)
S85
Evaluating the DevOps Reference Architecture for Multi-cloud IoT-Applications
SN Computer Science
2021 (Bou Ghantous and Gill, 2021)
S86
Hybrid IT and Multi Cloud an Emerging Trend and Improved Performance in Cloud Computing
SN Computer Science
2020 (Gundu et al., 2020)